### v1

# Ill-posedness of the Dirichlet problem for 2D Lagrangian mean curvature equation

Abstract.
We investigate the Dirichlet problem of the two dimensional Lagrangian mean curvature equation in a bounded domain. Infinitely many $C^{1,\alpha}(\alpha\in(0,\frac{1}{5}))$ very weak solutions are built through Nash-Kuiper construction. Moreover, we note there are infinitely many $C^{1,\alpha}$ very weak solutions that can not be improved to be $C^{2,\alpha}$.
1. Introduction
In the present paper, we study the Dirichlet problem for two dimensional Lagrangian mean curvature equation
(1.1) $\begin{cases}\cos\Theta\Delta v+\sin\Theta\left(\det\nabla^{2}v-1\right)=0\text{ in }\Omega,\\ v=g\text{ on }\partial\Omega,\end{cases}$
where $\Omega\subset\operatorname{\mathbb{R}}^{2}$ is a bounded domain, phase function $\Theta:\overline{\Omega}\to(-\pi,\pi)$ is a given $C^{2,\kappa}(\kappa\in(0,1))$ function. The equation
(1.2) $\cos\Theta\Delta v+\sin\Theta\left(\det\nabla^{2}v-1\right)=0$
has an equivalent form
(1.3) $\arctan\lambda_{1}+\arctan\lambda_{2}=\Theta,$
with $\lambda_{1},\lambda_{2}$ being eigenvalues of $\nabla^{2}v$. The phase function $\Theta$ is a potential for the mean curvature of the Lagrangian submanifold
$\Sigma:=\{(x,\nabla v(x));x\in\Omega\}$
or called the gradient graph.
Generally, $n$-dimensional Lagrangian mean curvature equation is
(1.4) $\sum_{k=1}^{n}\arctan\lambda_{k}=\Theta,$
where $\lambda_{i}$’s are eigenvalues of $\nabla^{2}v$. The mean curvature vector of $n$-dimensional gradient graph $\Sigma$ is equal to $J\nabla_{g}\Theta$, where $g=\mathrm{Id}+(\nabla^{2}v)^{2}$ is the induced metric on $\Sigma$, and $J(x,y)=(-y,x)$ is a rotation in $\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n}$. In particular, when $\Theta$ is a constant, (1.4) is called the special Lagrangian equation, which originates in the special Lagrangian geometry introduced by Harvey-Lawson [14]. Note $\Sigma$ is a minimal surface in $\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n}$ if and only if $\Theta$ is a constant.
1.1. Background
[…]
In this paper, we consider not only classical solutions but also very weak solutions to (1.1). In fact, very weak solution to the Monge-Ampère equation was first introduced by Lewicka-Pakzad [18] via very weak Hessian defined by Iwaniec [16]. Two dimensional very weak Hessian coincides with the Monge-Ampère measure for any convex solution ([11]). Moreover, Without boundary conditions, Lewicka-Pakzad [18], Székelyhidi and the first author [7], and Hirsch-Inauen and the first author [5] proved that there are infinitely many very weak solutions to the two dimensional Monge-Ampère equation of different regularity. After imposing boundary data, the first author in [4] studied very weak solutions to the Monge-Ampère equation. Recently, Li-Qiu [19] also constructed infinitely many very weak solutions of the Dirichlet problem for 2-Hessian equation. All the above approaches to construct very weak solutions rely on the Nash-Kuiper construction, which is originated from isometric embeddings (Nash-Kuiper theorem, see [22, 17, 9, 6, 10, 23]), and developed to be convex integration to prove $h$-principle (see [13]). In recent decades, the technique of convex integration become a powerful approach to show non-uniqueness of nonlinear partial differential equations.
1.2. Main result and strategies
Along the line of very weak solutions to the Monge-Ampère equation, we also introduce the concept of very weak solution to the Dirichlet problem (1.1).
Definition 1.1 (Very weak solution).
$v\in C^{1}(\Omega)$ is a very weak solution to the Dirichlet problem (1.1) in $\Omega$, if $v=g$ on $\partial\Omega$, and it solves the following equation in the very weak sense in $\Omega$,
(1.5) $\cos\Theta\mbox{ curl curl }(v\cdot\mbox{Id})+\sin\Theta\left(-\frac{1}{2}\mbox{ curl curl }(\nabla v\otimes\nabla v)-1\right)=0,$
namely, for any $\phi\in C^{\infty}_{0}(\Omega)$
$\int_{\Omega}\partial_{1}v\partial_{2}v\partial_{12}^{2}(\phi\sin\Theta)-\frac{1}{2}(\partial_{1}v)^{2}\partial_{22}^{2}(\phi\sin\Theta)-\frac{1}{2}(\partial_{2}v)^{2}\partial_{11}^{2}(\phi\sin\Theta)dx$
[…]
Now we are ready to state our main result.
Theorem 1.1.
Let $\Omega\subset\operatorname{\mathbb{R}}^{2}$ be a bounded simply connected domain with $C^{2,\kappa}$ boundary $\partial\Omega$ for any given $\kappa\in(0,1)$. Suppose $g\in C^{2,\kappa}(\partial\Omega)$ and the phase $\Theta:\overline{\Omega}\to(-\pi,\pi)$ is a $C^{2,\kappa}$ function. Then the following statements hold.
  • (1)
    If $|\sin\Theta|\leq c_{1}$ for some small positive constant $c_{1}$, then the Dirichlet problem (1.1) admits a unique $C^{2,\kappa}(\Omega)$ classical solution.
  • (2)
    If $|\sin\Theta|\geq c_{2}$ for any positive constant $c_{2}\in(0,1]$, then the Dirichlet problem (1.1) admits infinitely many $C^{1,\alpha}(\Omega)$ very weak solutions for any $\alpha\in(0,\frac{1}{5})$.
Remark 1.1.
The assumption of simply connected is just from the application of Lemma A.2 in the proof of Proposition 3.1.
Note $c_{2}$ is arbitrary, as a quick observation, we have
Corollary 1.2.
For any $c_{3}\in(0,c_{1})$, if $c_{3}\leq|\sin\Theta|\leq c_{1}$, then there are infinitely many $C^{1,\alpha}$ very weak solutions to the Dirichlet problem (1.1) which can not be improved to be $C^{2,\alpha}$.
Our strategy to show Theorem 1.1 (1) is considering (1.2) as a perturbation of Laplacian equation and then applying the implicit function theorem. As for Theorem 1.1 (2), with $|\sin\Theta|\geq c_{2}>0$, the problem is reduced to solve
$\cot\Theta\Delta v+\det\nabla^{2}v-1=0\text{ in the very weak sense}.$
We will solve the above equation by the Nash-Kuiper construction. The main obstacle is how to treat terms with phase function, comparing with [4], which will be regarded as a Laplacian of some function, namely,
$\cot\Theta\Delta v=\Delta(v\cot\Theta-V),\text{ in the very weak sense,}$,
with
$\Delta V=2\nabla v\cdot\nabla\cot\Theta+v\Delta\cot\Theta.$
Hence, the problem is turned to solve
curlcurl( sym $\mathrm{curl}\,\mathrm{curl}\,\big{(}\mbox{ sym }\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id}\big{)}=-1,$
in the very weak sense, and then $V\cdot\textrm{Id}$ becomes an error term of lower order. With the elliptic estimate in Lemma 3.4, we are able to control it and arrive at our conclusion.
We remark that the method proving Theorem 1.1 (2) has more general application.
Remark 1.2.
Let
$F_{\tau}(\lambda)=\left\{\begin{aligned} &\frac{1}{2}\big{(}\ln\lambda_{1}+\ln\lambda_{2}\big{)},\qquad\tau=0,\\ &\frac{\sqrt{a^{2}+1}}{2b}\big{(}\ln\frac{\lambda_{1}+a-b}{\lambda_{1}+a+b}+\ln\frac{\lambda_{2}+a-b}{\lambda_{2}+a+b}\big{)},\qquad 0<\tau<\frac{\pi}{4},\\ &-\sqrt{2}\big{(}\frac{1}{1+\lambda_{1}}+\frac{1}{1+\lambda_{2}}\big{)},\qquad\tau=\frac{\pi}{4},\\ &\frac{\sqrt{a^{2}+1}}{b}\big{(}\arctan\frac{\lambda_{1}+a-b}{\lambda_{1}+a+b}+\arctan\frac{\lambda_{2}+a-b}{\lambda_{2}+a+b}\big{)},\qquad\frac{\pi}{4}<\tau<\frac{\pi}{2},\\ &\arctan\lambda_{1}+\arctan\lambda_{2},\qquad\tau=\frac{\pi}{2},\end{aligned}\right.$
where $a=\cot\tau$, $b=\sqrt{|\cot^{2}\tau-1|}$. With slight modifications, we can establish result like Theorem 1.1 (2) to the following Dirichlet problem
(1.7) $\begin{cases}F_{\tau}(\lambda)=\Theta,&\mbox{ in }\Omega,\\ \qquad v=g,&\mbox{ on }\partial\Omega,\end{cases}$
where $\lambda_{1}$, $\lambda_{2}$ are eigenvalues of $\nabla^{2}v$. We note the equations $F_{\tau}(\lambda)=\Theta$ can be formulated to equations of the following form
$P\Delta u+Q\det\nabla^{2}u=R.$
Equations in (1.7) are originated from the study of gradient graphs $\Sigma$. By Warren [26], if $\Sigma$ is a minimal Lagrangian graph in $(\operatorname{\mathbb{R}}^{2}\times\operatorname{\mathbb{R}}^{2},g_{\tau})$ with
$g_{\tau}=\sin\tau\big{(}\sum_{i=1}^{2}\mathrm{d}x_{i}\otimes\mathrm{d}x_{i}+\sum_{j=1}^{2}\mathrm{d}y_{j}\otimes\mathrm{d}y_{j}\big{)}+\cos\tau\big{(}\sum_{i=1}^{2}\mathrm{d}x_{i}\otimes\mathrm{d}y_{i}+\sum_{j=1}^{2}\mathrm{d}y_{j}\otimes\mathrm{d}x_{j}\big{)},$
then
$F_{\tau}(\lambda(\nabla^{2}v))=\text{constant},$
We refer readers to results about gradient graphs in general dimensions in [26].
1.3. Organization and notations
The rest of the paper is organized as follows. Theorem 1.1 (1) will be proved in Section 2. Our stage proposition, which is our main contribution, and the proof for Theorem 1.1 (2) will be presented in Section 3 and Section 4 respectively.
Finally, we provide some notations for convenience. For any function $h:\Omega\to\operatorname{\mathbb{R}}$, set
$\|h\|_{0}=\sup_{\Omega}|h|,\,\,\|h\|_{m}=\sum_{j=0}^{m}\max_{|\beta|=j}\|\partial^{\beta}h\|_{0},$
$\|h\|_{m+\alpha}=\|h\|_{m}+[h]_{m+\alpha},$
where the Hölder semi-norm is defined by
$[h]_{m+\alpha}=\max_{|\beta|=m}\sup_{x\neq y}\frac{|\partial^{\beta}h(x)-\partial^{\beta}h(y)|}{|x-y|^{\alpha}}.$
Here $m$ is a nonnegative integer, $\alpha\in(0,1]$ a constant, and $\beta$ be a multi-index. We denote by $B_{r}$ the ball in $\operatorname{\mathbb{R}}^{2}$ with radius $r$ centering at 00. Let $\varphi\in C^{\infty}_{0}(B_{1})$ be a standard two dimensional mollifier. We define $\varphi_{l}(x)=l^{-2}\varphi(l^{-1}x)$, and $l>0$ is called the length-scale. We also denote
[…]
Some properties of mollification can be found in Lemma A.1. The operator $\mathrm{curl}$ in two dimension is defined by
[…]
Besides, for $\xi=(\xi_{1},\xi_{2})\in\mathbb{R}^{2}$ and $\eta=(\eta_{1},\eta_{2})\in\mathbb{R}^{2}$, we write $\xi\otimes\eta$ a $2\times 2$-matrix with entries $(\xi\otimes\eta)_{ij}=\xi_{i}\eta_{j}$. By a direct calculation, for a $C^{3}$-function $w$, we have
$\det\nabla^{2}w=-\frac{1}{2}\mbox{curl curl }(\nabla w\otimes\nabla w)\mbox{ and }\Delta w=\mbox{curl curl }(w\cdot\mathrm{Id}).$
For a $C^{1}$-function $w:\Omega\to\operatorname{\mathbb{R}}^{2}$, note $\nabla w=(\partial_{i}w_{j})_{2\times 2}$ and
$\mathrm{sym}\nabla w=\frac{1}{2}(\nabla w+\nabla w^{T}).$
Acknowledgement
Wentao Cao’s research is supported in part by the National Natural Science Foundation of China (Grant No. 12471224). Zhehui Wang’s research is supported in part by Young Scientists Fund of the National Natural Science Foundation of China (Grant No. 12401247), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515110910).
2. Proof of Theorem 1.1 (1)
The proof of Theorem 1.1 is based on an application of the implicit function theorem.
Proof.
Let $\widetilde{g}\in C^{2,\kappa}(\overline{\Omega})$ such that $\widetilde{g}=g$ on $\partial\Omega$. Set
$\mathcal{X}=\{\psi\in C^{2,\kappa}(\overline{\Omega}):\psi=0\text{ on }\partial\Omega\}.$
Define $G:\mathcal{X}\times C^{\kappa}(\overline{\Omega})\rightarrow C^{\kappa}(\overline{\Omega})$ by
$G(\psi,\Psi)=\Delta\psi+\Psi\cdot(\det\nabla^{2}(\psi+\widetilde{g})-1)+\Delta\widetilde{g}.$
Then we have the following conclusion.
Claim: There is a unique $\psi\in\mathcal{X}$ and a small enough constant $\mu\in(0,1)$ such that $G(\psi,\Psi)=0$ for any $\Psi$ satisfying $\|\Psi\|_{\kappa}\leq\mu$.
In fact, the Fréchet derivative of $G$ with respect to $\psi$ is given by
$G_{\psi}(\psi,\Psi)\vartheta=\Delta\vartheta+\frac{d\,}{dt}\Big{|}_{t=0}\Psi\cdot(\det\nabla^{2}(\psi+\widetilde{g}+t\vartheta)-1),$
and hence $G_{\psi}(\psi,0)\vartheta=\Delta\vartheta$ for any $\vartheta\in\mathcal{X}$. Therefore, for any $\psi\in\mathcal{X}$, $G_{\psi}(\psi,0)$ is an invertible operator from $\mathcal{X}$ to $C^{\kappa}(\overline{\Omega})$.
Next, take $\psi_{0}\in\mathcal{X}$ such that $\Delta\psi_{0}=-\Delta\widetilde{g}$. The existence of such $\psi_{0}$ is provided by the Schauder theory. By the implicit function theorem, the equation $G(\psi,\Psi)=0$ admits a unique solution $\psi=\psi_{\Psi}\in\mathcal{X}$ for each $\Psi$ with $\|\Psi\|_{C^{\kappa}(\overline{\Omega})}\leq\mu$, where $\mu$ is a sufficiently small constant. Moreover, $\psi_{\Psi}$ is close to $\psi_{0}$ in the $C^{2,\kappa}(\overline{\Omega})$-norm.
Let $\Theta:\overline{\Omega}\to(-\pi,\pi)$ be a $C^{1,\kappa}$ function. By the interpolation, we note
$\|f\|_{C^{\kappa}(B_{r})}\leq\sigma\|f\|_{C^{1,\kappa}(B_{r})}+(\frac{C}{\sigma^{\kappa}}+1)\|f\|_{L^{\infty}(B_{r})}$
for any $0<\sigma\leq r$ and $f\in C^{1,\kappa}(B_{r})$. Hence, if $|\sin\Theta|\leq c_{1}$ with small constant $c_{1}$ to be fixed, then we can get
$\|\sin\Theta\|_{\kappa}\leq C(\kappa,\|\Theta\|_{1+\kappa})(c_{1}+c_{1}^{1-\kappa}),$
where $C(\kappa,\|\Theta\|_{1+\kappa})$ is a constant depending only on $\kappa,\|\Theta\|_{1+\kappa}.$ By a direct computation,
$\|\tan\Theta\|_{\kappa}$ $\leq\|\sin\Theta\|_{0}\|\frac{1}{\cos\Theta}\|_{\kappa}+\|\sin\Theta\|_{\kappa}\|\frac{1}{\cos\Theta}\|_{0}$
$\leq 4\|\sin\Theta\|_{\kappa}\|\frac{1}{\cos\Theta}\|_{1}$
$\leq C(\kappa,\|\Theta\|_{1+\kappa})\|\sin\Theta\|_{\kappa}(\|\frac{1}{\cos\Theta}\|_{0}+\|\frac{1}{\cos^{2}\Theta}\|_{0}\|\sin\Theta\|_{0})$
$\leq C(\kappa,\|\Theta\|_{1+\kappa})(c_{1}+c_{1}^{1-\kappa})(\frac{1}{\sqrt{1-c_{1}^{2}}}+\frac{c_{1}}{1-c_{1}^{2}}).$
Now, we choose $c_{1}$ small enough such that $\|\tan\Theta\|_{C^{\kappa}(\overline{\Omega})}\leq\mu$. By the above Claim, there is a unique $\psi\in\mathcal{X}$ such that
$\Delta\psi+\tan\Theta(\det\nabla^{2}(\psi+\widetilde{g})-1)+\Delta\widetilde{g}=0.$
Then, $v:=\psi+\widetilde{g}$ uniquely solves
$&\Delta v+\tan\Theta(\det\nabla^{2}v-1)=0\mbox{ in }\Omega,\\ &v=g\mbox{ on }\partial\Omega.\end{aligned}\right.$
Hence, we complete the proof of Theorem 1.1 (1). ∎
3. Stage Proposition
The rest of the paper is devoted to show Theorem 1.1 (2). Similar to the treatment in [4, 7, 18] for the Monge-Ampère equation, we split (1.2) (after dividing $\sin\Theta$) into a differential system
(3.1) curl curl $-\mbox{curl curl }A=1,$
(3.2) $A=\mbox{ sym }\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id},$
where $V$ is the solution to
$\Delta V&=F(v),\mbox{ in }\Omega,\\ V&=0,\mbox{ on }\partial\Omega,\end{aligned}\right.$
and $F(v)$ is defined as
(3.3) $F(v)=2\nabla v\cdot\nabla\cot\Theta+v\Delta\cot\Theta.$
In fact, by a direct calculation, we note that the $v$ satisfying (3.1)-(3.2) solves (1.5) in the very weak sense.
To solve (3.2), our strategy is applying the technique of Nash-Kuiper construction to construct a sequence of subsolutions satisfying boundary condition, whose limit is our desired solution. Set
(3.4) $D(v,w):=\mbox{ sym }\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id},$
then such subsolutions, which are called adapted subsolutions, are defined as below.
Definition 3.1.
A pair of mappings $(v,w):\overline{\Omega}\to\operatorname{\mathbb{R}}\times\operatorname{\mathbb{R}}^{2}$ is called a $C^{1,\beta}(\beta<1)$ adapted subsolution of (3.2) if $(v,w)\in C^{1,\beta}$ and
$A-D(v,w)=\rho^{2}(\textrm{Id}+H)\mbox{ in }\overline{\Omega},$
where $D(v,w)$ is as (3.4), and $\rho\in C(\overline{\Omega})$ is a nonnegative function with $\partial\Omega\subset\{x:\rho(x)=0\}$, and $H\in C(\overline{\Omega};\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ is a symmetric matrix such that $\textrm{Id}+H$ is positive definite.
Note that the definition is a slight modification of that in [4]. To construct new adapted subsolutions, we need to add the deficit $A-D(v,w)$ to the given adapted subsolution $v.$ Since the deficit is of the form $\rho^{2}(\mathrm{Id}+H)$, we require the following stage proposition (in the terminology of Nash [22]) for iteration, which is parallel to Proposition 3.1 in [4] and whose proof is different due to the new error terms. The new error terms are then controlled through Lemma 3.29.
Proposition 3.1.
Suppose $\Sigma\subset\operatorname{\mathbb{R}}^{2}$ is a simply connected open bounded set with $C^{2}$ boundary. Let $v\in C^{2}(\overline{\Sigma})$ and $w\in C^{2}(\Sigma,\operatorname{\mathbb{R}}^{2})$ such that
(3.5) $\|v\|_{1}+\|w\|_{1}\leq\gamma,$
(3.6) $\|v\|_{2}+\|w\|_{2}\leq\delta^{\frac{1}{2}}\lambda,$
and let $\rho\in C^{1}(\overline{\Sigma})$, $H\in C^{1}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ such that
(3.7) $\|\rho\|_{0}\leq\delta^{\frac{1}{2}},\quad\|\nabla\rho\|_{0}\leq\delta^{\frac{1}{2}}\lambda,$
(3.8) $\|H\|_{0}\leq\lambda^{-\alpha},\quad\|\nabla H\|_{0}\leq\lambda^{1-\alpha},$
for some $\gamma>1$, $10<\delta,\,\alpha<$ 1 and $\lambda>1$ with $2\leq\lambda^{\alpha}$ and $\delta^{1/2}\lambda>1$. Then for any $\tau>1$, there exist functions $v^{\star}\in C^{2}(\overline{\Sigma})$, $w^{\star}\in C^{2}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2})$ and symmetric matrix $\operatorname{\mathcal{E}}\in C^{1}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ such that
(3.9) $D(v^{\star},w^{\star})=D(v,w)+\rho^{2}(\mathrm{Id}+H)+\operatorname{\mathcal{E}},\,\,\mbox{in }\overline{\Sigma},$
(3.10) $(v^{\star},w^{\star})=(v,w),\,\,\operatorname{\mathcal{E}}=0,\,\,\mbox{on }\overline{\Sigma}\setminus(\mathrm{supp}\rho+B_{\lambda^{-\tau}}),$
and they satisfy the following estimates
(3.11) $\|v^{\star}-v\|_{0}+\|w^{\star}-w\|_{0}$ $\leq C\delta^{\frac{1}{2}}\lambda^{-\tau},$
(3.12) $\|v^{\star}-v\|_{1}+\|w^{\star}-w\|_{1}$ $\leq C\delta^{\frac{1}{2}},$
(3.13) $\|v^{\star}\|_{2}+\|w^{\star}\|_{2}$ $\leq C\delta^{\frac{1}{2}}\lambda^{2\tau-1},$
and
(3.14) $\|\operatorname{\mathcal{E}}\|_{0}\leq\bar{C}\delta\lambda^{1-\tau},\quad\|\operatorname{\mathcal{E}}\|_{1}\leq\bar{C}\delta\lambda^{\tau},$
where constant $C$ and $\bar{C}$ depend only on $\gamma$ and $\alpha$.
Comparing the proof of Proposition 3.1 in [4], the key ingredient and difference of the proof for our Proposition 3.1 is applying the elliptic estimate in Lemma 3.4 to control the new error terms with the phase function.
Proof.
Same as [4], we also divide the proof into three steps.
Step 1. Mollification. Take the lengthscale $\ell=\lambda^{-\tau}$, Set $\tilde{\rho}=\rho*\varphi_{\ell}$ and $\tilde{H}=H*\varphi_{\ell}$. Then by (3.7), (3.8) and Lemma A.1, we know,
$\|\tilde{\rho}\|_{0}\leq C\delta^{\frac{1}{2}},\quad\|\tilde{\rho}\|_{j}\leq C(j)\ell^{1-j}[\rho]_{1}\leq C(j)\delta^{\frac{1}{2}}\lambda\ell^{1-j},$
$\|\tilde{\rho}-\rho\|_{0}\leq C\ell[\rho]_{1}\leq C\delta^{\frac{1}{2}}\lambda\ell,$
$\|\tilde{H}\|_{0}\leq C\lambda^{-\alpha},\quad\|\tilde{H}\|_{j}\leq C(j)\ell^{1-j}[H]_{1}\leq C(j)\lambda^{1-\alpha}\ell^{1-j},$
$\|\tilde{H}-H\|_{0}\leq C\ell[H]_{1}\leq C\lambda^{1-\alpha}\ell,$
for $j\in\operatorname{\mathbb{N}}_{+}$. Let $\tilde{h}=\tilde{\rho}^{2}(\mathrm{Id}+\tilde{H})$ and $h=\rho^{2}(\mathrm{Id}+{H})$, then
$\|\tilde{h}-h\|_{0}$ $\leq\|\tilde{\rho}^{2}-\rho^{2}\|_{0}+\|\tilde{\rho}^{2}\tilde{H}-\rho^{2}H\|_{0}\leq C\delta\lambda\ell+C\delta\lambda^{1-\alpha}\ell\leq C\delta\lambda^{1-\tau},$
$\|\tilde{h}-h\|_{1}$ $\leq\|\tilde{\rho}^{2}-\rho^{2}\|_{1}+\|\tilde{\rho}^{2}(\tilde{H}-H)\|_{1}+\|(\tilde{\rho}^{2}-\rho^{2})H\|_{1}$
$\leq C(\delta\lambda+\delta\lambda^{1-\alpha})\leq C\delta\lambda.$
Besides, it holds that
$+\tilde{H}\geq(1-\lambda^{-\alpha})\textrm{Id}>\tfrac{1}{2}\textrm{Id},$ Id ,
and
$\|\textrm{Id}+\tilde{H}\|_{\alpha^{\prime}}\leq 1+\|\tilde{H}\|_{\alpha^{\prime}}\leq C(1+\lambda^{\alpha^{\prime}-\alpha})\leq C,$
provided $0<\alpha^{\prime}<\alpha.$ Thus applying Lemma A.2 to $\textrm{Id}+\tilde{H}$, we get
$\textrm{Id}+\tilde{H}=\tilde{a}^{2}(\nabla\Phi_{1}\otimes\nabla\Phi_{1}+\nabla\Phi_{2}\otimes\nabla\Phi_{2}),$
with $\tilde{a},\Phi=(\Phi_{1},\Phi_{2})$ satisfying
$\begin{split}&\tilde{a}\geq C^{-1},\quad\det(\nabla\Phi)\geq C^{-1},\\ &\|\tilde{a}\|_{j+\alpha^{\prime}}+\|\nabla\Phi\|_{j+\alpha^{\prime}}\leq\|\mathrm{Id}+\tilde{H}\|_{j+\alpha^{\prime}}\\ \leq&C(1+\|\tilde{H}\|_{j}\ell^{-\alpha^{\prime}})\leq C(j,\alpha^{\prime},\alpha)\lambda^{1-\alpha}\ell^{1-j-\alpha^{\prime}},\end{split}$
for any $j\geq 1$, and
$\|\tilde{a}\|_{\alpha^{\prime}}+\|\nabla\Phi\|_{\alpha^{\prime}}\leq\|\mathrm{Id}+\tilde{H}\|_{\alpha^{\prime}}\leq C,$
Taking $\alpha^{\prime}\leq\frac{\alpha}{\tau},$ we have
(3.15) $C^{-1}\leq|\nabla\Phi_{1}|,\,|\nabla\Phi_{2}|\leq C,\quad\|\nabla\Phi\|_{j}\leq C(j)\lambda\ell^{1-j}\leq C(j)\lambda^{1-\tau+\tau j},$
and
$\|\tilde{a}\|_{j}\leq C(j)\lambda^{1-\tau+\tau j}\text{ for any }j\geq 1.$
Setting $a=\tilde{a}\tilde{\rho}$, we obtain
$\tilde{h}=a^{2}(\nabla\Phi_{1}\otimes\nabla\Phi_{1}+\nabla\Phi_{2}\otimes\nabla\Phi_{2}),$
and
(3.16) $\|a\|_{0}\leq C\delta^{1/2},\quad\|a\|_{j}\leq C(j)\delta^{1/2}\lambda^{1-\tau+\tau j}$
for any $j\geq 1.$
Step 2. Adding the first rank-one matrix. Set $\tilde{v}=v*\varphi_{\ell}$ and $\tilde{w}=w*\varphi_{\ell}$. With Lemma A.1, we get
(3.17) $\|\tilde{v}-v\|_{1}+\|\tilde{w}-w\|_{1}\leq C\ell(\|v\|_{2}+\|w\|_{2})\leq C\delta^{\frac{1}{2}}\lambda^{1-\tau},$
(3.18) $\|\tilde{v}\|_{2+j}+\|\tilde{w}\|_{2+j}\leq C\ell^{-j}(\|v\|_{2}+\|w\|_{2})\leq C\delta^{\frac{1}{2}}\lambda^{1+\tau j},$
Define
$v^{\lozenge}(x)=$ $v(x)+\lambda^{-\tau}\Gamma_{1}(a(x),\lambda^{\tau}\Phi_{1}(x)),$
$w^{\lozenge}(x)=$ $w(x)-\lambda^{-\tau}\Gamma_{1}(a(x),\lambda^{\tau}\Phi_{1}(x))\nabla\tilde{v}(x)$
$+\lambda^{-\tau}\Gamma_{2}(a(x),\lambda^{\tau}\Phi_{1}(x))\nabla\Phi_{1}(x),$
where $\Gamma_{i},i=1,2$ are defined as follows.
(3.19) $\Gamma_{1}(s,t)=\frac{s}{\pi}\sin(2\pi t),\quad\Gamma_{2}(s,t)=-\frac{s^{2}}{4\pi}\sin(4\pi t).$
It is easy to see that $\Gamma_{i}(s,t)=\Gamma_{i}(s,t+1)$, $i=1,2$ and
(3.20) $\frac{1}{2}|\partial_{t}\Gamma_{1}(s,t)|^{2}+\partial_{t}\Gamma_{2}(s,t)=s^{2}.$
From both definitions of $\Gamma_{i},i=1,2$ we gain
(3.21) $(v^{\lozenge},w^{\lozenge})=(v,w)\textit{ on }\overline{\Omega}\setminus\textrm{supp }a.$
Observe that
(3.22) $\|v^{\lozenge}-v\|_{j}\leq\lambda^{-\tau}\|\Gamma_{1}\|_{j},$
and
(3.23) $\begin{split}\|w^{\lozenge}-w\|_{j}\leq C(j)\lambda^{-\tau}(&\|\Gamma_{1}\|_{j}\|\tilde{v}\|_{1}+\|\Gamma_{1}\|_{0}\|\tilde{v}\|_{j+1}\\ &+\|\Gamma_{2}\|_{j}\|\nabla\Phi_{1}\|_{0}+\|\Gamma_{2}\|_{0}\|\nabla\Phi_{1}\|_{j}),\end{split}$
for $j=0,1,2.$ Here $\|\Gamma_{i}\|_{j}$ denotes $C^{j}$ norms of function $x\rightarrow\Gamma_{i}(a(x),\lambda^{\tau}\Phi_{1}(x)).$
Lemma 3.2.
Assume all the assumptions of Proposition 3.1 are fulfilled, then the following estimates hold.
(3.24) $\begin{split}&\|\Gamma_{1}\|_{0}+\|\partial_{t}\Gamma_{1}\|_{0}+\|\partial_{t}^{2}\Gamma_{1}\|_{0}\leq C\delta^{1/2},\\ &\|\Gamma_{1}\|_{1}+\|\partial_{t}\Gamma_{1}\|_{1}\leq C\delta^{1/2}\lambda^{\tau},\\ &\|\partial_{s}\Gamma_{1}\|_{j}\leq C\lambda^{\tau j},\,j=0,1,\\ &\|\Gamma_{1}\|_{2}\leq C\delta^{1/2}\lambda^{2\tau},\end{split}$
and
(3.25) $\begin{split}&\|\Gamma_{2}\|_{0}+\|\partial_{t}\Gamma_{2}\|+\|\partial_{t}^{2}\Gamma_{2}\|_{0}\leq C\delta,\\ &\|\Gamma_{2}\|_{1}+\|\partial_{t}\Gamma_{2}\|_{1}\leq C\delta\lambda^{\tau},\\ &\|\partial_{s}\Gamma_{2}\|_{j}\leq C\delta^{1/2}\lambda^{\tau j},\,j=0,1,\\ &\|\Gamma_{2}\|_{2}\leq C\delta\lambda^{2\tau}.\end{split}$
The proof of Lemma 3.2 is same as that in [4], but for the completeness, we put it in Appendix.
So from (3.22), (3.15), (3.23), and Lemma 3.2, we get
(3.26) $\|v^{\lozenge}-v\|_{j}\leq C\delta^{1/2}\lambda^{\tau j-\tau},\,\,j=0,1,2.$
and
(3.27) $\|w^{\lozenge}-w\|_{j}\leq C(\delta^{1/2}\lambda^{\tau j-\tau}+\delta\lambda^{\tau j-\tau})\leq C\delta^{1/2}\lambda^{\tau j-\tau},j=0,1,2.$
Hence by (3.5), (3.6), (3.26) and (3.27), one arrives at
(3.28) $\begin{split}&\|v^{\lozenge}\|_{1}+\|w^{\lozenge}\|_{1}\leq C,\\ &\|v^{\lozenge}\|_{2}+\|w^{\lozenge}\|_{2}\leq C\delta^{1/2}\lambda^{\tau}.\end{split}$
Next we shall control $C^{j},j=0,1$ norms of the error matrix $\mathcal{E}.$ Taking derivatives gives
$\nabla v^{\lozenge}$ $=\nabla v+\lambda^{-\tau}\partial_{s}\Gamma_{1}\nabla a+\partial_{t}\Gamma_{1}\nabla\Phi_{1},$
$\nabla w^{\lozenge}$ $=\nabla w-\lambda^{-\tau}\partial_{s}\Gamma_{1}\nabla\tilde{v}\otimes\nabla a-\partial_{t}\Gamma_{1}\nabla\tilde{v}\otimes\nabla\Phi_{1}-\lambda^{-\tau}\Gamma_{1}\nabla^{2}\tilde{v}$
$\quad+\lambda^{-\tau}\partial_{s}\Gamma_{2}\nabla\Phi_{1}\otimes\nabla a+\partial_{t}\Gamma_{2}\nabla\Phi_{1}\otimes\nabla\Phi_{1}.$
Hence, with (3.20), we have
$\mathcal{E}_{1}=$ $D(v^{\lozenge},w^{\lozenge})-(D(v,w)+a^{2}\nabla\Phi_{1}\otimes\nabla\Phi_{1})=\sum_{k=1}^{5}\mathcal{E}_{1k}$
with
$\mathcal{E}_{11}:=$ $\lambda^{-\tau}(\partial_{t}\Gamma_{1}\partial_{s}\Gamma_{1}\mathrm{sym}(\nabla\Phi_{1}\otimes\nabla a)-\Gamma_{1}\nabla^{2}\tilde{v}+\partial_{s}\Gamma_{2}\textrm{sym}(\nabla\Phi_{1}\otimes\nabla a)),$
$\mathcal{E}_{12}:=$ ${\lambda}^{-\tau}\partial_{s}\Gamma_{1}\textrm{sym}((\nabla v-\nabla\tilde{v})\otimes\nabla a)+\partial_{t}\Gamma_{1}\textrm{sym}((\nabla v-\nabla\tilde{v})\otimes\nabla\Phi_{1}),$
$\mathcal{E}_{13}:=$ $\tfrac{1}{2}\lambda^{-2\tau}(\partial_{s}\Gamma_{1})^{2}\nabla a\otimes\nabla a,$
$\mathcal{E}_{14}:=$ $(V^{\lozenge}-V)\cdot\mathrm{Id},\,\,(\text{ here }\Delta V^{\lozenge}=F(v^{\lozenge}))$
$\mathcal{E}_{15}:=$ $((v^{\lozenge}-v)\cot\Theta)\cdot\mathrm{Id}$
Estimates of the first three errors and their proof are same as those in [4].
Lemma 3.3.
Assume all the assumptions of Proposition 3.1 are fulfilled, then for $j=0,1$, the following estimates hold.
$\|\mathcal{E}_{11}\|_{j}\leq C\delta\lambda^{1-\tau(1-j)},$
$\|\mathcal{E}_{12}\|_{j}\leq C\delta\lambda^{1-\tau(1-j)},$
$\|\mathcal{E}_{13}\|_{j}\leq C\delta\lambda^{2-j-2\tau(1-j)}.$
For completeness, we put the proof of Lemma 3.3 in Appendix.
To estimate $\operatorname{\mathcal{E}}_{14}$, we set
$\zeta=2(v^{\lozenge}-v)\nabla\cot\Theta,\quad f=-(v^{\lozenge}-v)\Delta\cot\Theta.$
Note that
$\Delta(V^{\lozenge}-V)=F(v^{\lozenge})-F(v)&=\operatorname{div}\zeta+f,\mbox{ in }\Omega,\\ V^{\lozenge}-V&=0,\mbox{ on }\partial\Omega.\end{aligned}\right.$
Hence we require the follow elliptic estimates.
Lemma 3.4.
Let $\Omega\subset\operatorname{\mathbb{R}}^{n}$ be a bounded domain. Suppose $\zeta=(\zeta_{1},\cdots,\zeta_{n})$ where $\zeta_{i}\in C^{1}(\Omega)\cap C^{0}(\overline{\Omega})$, $i=1,\cdots,n$, and $f\in C^{0}(\overline{\Omega})$. Let $u\in C^{2}(\Omega)\cap C^{0}(\overline{\Omega})$ be the solution to
$\begin{cases}\Delta u=\operatorname{div}\zeta+f,&\mbox{ in }\Omega,\\ \quad u=0,&\mbox{ on }\partial\Omega.\end{cases}$
Then,
(3.29) $\|u\|_{0}\leq C(n,\Omega)(\|f\|_{0}+\|\zeta\|_{0}),$
where $C(n,\Omega)$ is a constant depending only on $n$ and $\Omega$.
Proof.
Let $u_{1}$ be the solution of
$\begin{cases}\Delta u_{1}=\operatorname{div}\zeta,&\mbox{ in }\Omega,\\ \quad u_{1}=0,&\mbox{ on }\partial\Omega,\end{cases}$
and $u_{2}$ be the solution of
$\begin{cases}\Delta u_{2}=f,&\mbox{ in }\Omega,\\ \quad u_{2}=0,&\mbox{ on }\partial\Omega,\end{cases}$
then $u=u_{1}+u_{2}$, and
(3.30) $\|u\|_{0}\leq\|u_{1}\|_{0}+\|u_{2}\|_{0}.$
By the maximum principle,
(3.31) $\|u_{2}\|_{0}\leq C(n,\mathrm{diam}(\Omega))\|f\|_{0},$
where $\mathrm{diam}(\Omega)$ is the diameter of $\Omega$. Now (3.29) will follow by (3.30), (3.31) and the claim below:
Claim: $\|u_{1}\|_{0}\leq C(n,\Omega)\|\zeta\|_{0}.$
Let $\phi$ be the solution of
$\begin{cases}\Delta\phi=u_{1},&\mbox{ in }\Omega,\\ \quad\phi=0,&\mbox{ on }\partial\Omega,\end{cases}$
then
$\int_{\Omega}u_{1}^{2}\mathrm{d}x$ $=\int_{\Omega}u_{1}\Delta\phi\mathrm{d}x=\int_{\Omega}\phi\Delta u_{1}\mathrm{d}x=\int_{\Omega}\phi\operatorname{div}\zeta\mathrm{d}x$
$=-\int_{\Omega}\nabla\phi\cdot\zeta\mathrm{d}x\leq\|\nabla\phi\|_{L^{2}}\|\zeta\|_{L^{2}}.$
Besides, the maximum principle tells us that
(3.32) $\|\phi\|_{0}\leq C(n,\mathrm{diam}(\Omega))\|u_{1}\|_{0}.$
Note
$\int_{\Omega}|\nabla\phi|^{2}\mathrm{d}x=-\int_{\Omega}\phi\Delta\phi\mathrm{d}x=-\int_{\Omega}\phi u_{1}\mathrm{d}x\leq\|\phi\|_{L^{2}}\|u_{1}\|_{L^{2}},$
we have
$\|\nabla\phi\|_{L^{2}}\leq\|\phi\|_{L^{2}}^{\frac{1}{2}}\|u_{1}\|_{L^{2}}^{\frac{1}{2}},$
and further,
$\|u_{1}\|^{2}_{L^{2}}\leq\|\phi\|_{L^{2}}^{\frac{1}{2}}\|u_{1}\|_{L^{2}}^{\frac{1}{2}}\|\zeta\|_{L^{2}}.$
Therefore, we get
(3.33) $\|u_{1}\|^{\frac{3}{2}}_{L^{2}}\leq\|\phi\|_{L^{2}}^{\frac{1}{2}}\|\zeta\|_{L^{2}}\leq|\Omega|^{\frac{1}{4}}\|\phi\|_{0}^{\frac{1}{2}}|\Omega|^{\frac{1}{2}}\|\zeta\|_{0}\leq C(n,\Omega)\|u_{1}\|^{\frac{1}{2}}_{0}\|\zeta\|_{0}.$
here we apply (3.32) in last inequality. By Theorem 8.15 in [12], we know
$\|u_{1}\|_{0}\leq C(n,\Omega)(\|\zeta\|_{0}+\|u_{1}\|_{L^{2}}),$
which combined with (3.33) will imply
(3.34) $\|u_{1}\|_{0}\leq C(n,\Omega)(\|\zeta\|_{0}+\|u_{1}\|_{0}^{\frac{1}{3}}\|\zeta\|_{0}^{\frac{2}{3}}).$
Denote $D=\|u_{1}\|_{0}$ and $B=\|\zeta\|_{0}$. We consider the following two cases.
Case 1: $D^{\frac{1}{3}}B^{\frac{2}{3}}\leq B$. In this case, we have $D\leq B$ and the claim follows.
Case 2: $D^{\frac{1}{3}}B^{\frac{2}{3}}\geq B$. In this case, by (3.34) we have
$D\leq 2C(n,\Omega)D^{\frac{1}{3}}B^{\frac{2}{3}},$
which also implies
$D\leq C(n,\Omega)B.$
Then we complete the proof of the claim. ∎
With Lemma 3.4 and (3.26),
$\|V^{\lozenge}-V\|_{0}\leq C(\|f\|_{0}+\|\zeta\|_{0})\leq C\|v^{\lozenge}-v\|_{0}\leq C\delta^{\frac{1}{2}}\lambda^{-\tau},$
and for any $\sigma\in(0,1)$, by the $C^{1,\sigma}$ estimates ((4.45)-(4.46) in [12]) ,
$\|V^{\lozenge}-V\|_{1+\sigma}\leq C(\|\zeta\|_{\sigma}+\|f\|_{0})\leq$ $C\|v-v^{\lozenge}\|_{\sigma}\leq C\|v-v^{\lozenge}\|_{0}^{1-\sigma}\|v-v^{\lozenge}\|_{1}^{\sigma}$
$\leq$ $C(\delta^{\frac{1}{2}})^{\sigma}(\delta^{\frac{1}{2}}\lambda^{-\tau})^{1-\sigma}=C\delta^{\frac{1}{2}}\lambda^{-\tau(1-\sigma)}.$
Then,
$\|\mathcal{E}_{14}\|_{0}$ $=\|V-V^{\lozenge}\|_{0}\leq C\delta^{\frac{1}{2}}\lambda^{-\tau}\leq C\delta\lambda^{1-\tau},$
$\|\mathcal{E}_{14}\|_{1}$ $=\|V-V^{\lozenge}\|_{1}\leq\|V-V^{\lozenge}\|_{1+\sigma}\leq C\delta^{\frac{1}{2}}\lambda^{-\tau(1-\sigma)}\leq C\delta\lambda.$
It is remained to estimates $\mathcal{E}_{15}$. From (3.26), for $j=0,1$, we can derive
$\|\mathcal{E}_{15}\|_{j}\leq C\|v-v^{\lozenge}\|_{j}\leq C\delta^{\frac{1}{2}}\lambda^{\tau j-\tau}\leq\delta\lambda^{1-\tau(1-j)},$
where we have used $\delta^{1/2}\lambda>1.$ Putting the estimates of $\mathcal{E}_{1k},k=1,\cdots,5$ together, we obtain
(3.35) $\|\mathcal{E}_{1}\|_{0}\leq C\delta\lambda^{1-\tau},\quad\|\mathcal{E}_{1}\|_{1}\leq C\delta\lambda.$
Step 3.Adding the second rank-one matrix. We also mollify $(v^{\lozenge},w^{\lozenge})$ to get $(\tilde{v}^{\lozenge},\tilde{w}^{\lozenge})$ at length-scale $\lambda^{1-2\tau}$. Then by (3.28) and Lemma A.1, one also has
(3.36) $\begin{split}&\|\tilde{v}^{\lozenge}-v^{\lozenge}\|_{1}+\|\tilde{w}^{\lozenge}-w^{\lozenge}\|_{1}\leq C\delta^{1/2}\lambda^{1-\tau},\\ &\|\tilde{v}^{\lozenge}\|_{2+j}+\|\tilde{w}^{\lozenge}\|_{2+j}\leq C\delta^{1/2}\lambda^{\tau+(2\tau-1)j},\end{split}$
for any $j\geq 0.$ Set
$v^{\star}(x)=$ $v^{\lozenge}(x)+\lambda^{1-2\tau}\Gamma_{1}(a(x),\lambda^{2\tau-1}\Phi_{2}),$
$w^{\star}(x)=$ $w^{\lozenge}(x)-\lambda^{1-2\tau}\Gamma_{1}(a(x),\lambda^{2\tau-1}\Phi_{2})\nabla\tilde{v}^{\lozenge}(x)$
$+\lambda^{1-2\tau}\Gamma_{2}(a(x),\lambda^{2\tau-1}\Phi_{2})\nabla\Phi_{2}.$
By the definitions of $\Gamma_{i},i=1,2$, we get
$(v^{\star},w^{\star})=(v^{\lozenge},w^{\lozenge})\textit{ on }\overline{\Omega}\setminus\textit{supp }a.$
Then (3.10) follows from (3.21) and
dist(supp $\textit{dist}(\textrm{supp }a,\textrm{supp }\rho)=\lambda^{-\tau}.$
Furthermore, same as Step 2, using (3.16),(3.15) and (3.28), one is able to derive
(3.37) $\begin{split}\|v^{\star}-v^{\lozenge}\|_{0}\leq&\lambda^{1-2\tau}\|\Gamma_{1}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{0}\leq C\delta\lambda^{1-2\tau},\\ \|v^{\star}-v^{\lozenge}\|_{1}\leq&\lambda^{1-2\tau}\|\Gamma_{1}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{1}\leq C\delta,\\ \|v^{\star}-v^{\lozenge}\|_{2}\leq&\lambda^{1-2\tau}\|\Gamma_{1}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{2}\leq C\delta\lambda^{2\tau-1}.\end{split}$
and
(3.38) $\begin{split}\|w^{\star}-w^{\lozenge}\|_{j}\leq&C(j)\lambda^{1-2\tau}(\|\Gamma_{1}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{j}\|\tilde{v}^{\lozenge}\|_{1}\\ &\quad+\|\Gamma_{1}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{0}\|\tilde{v}^{\lozenge}\|_{j+1}\\ &\quad+\|\Gamma_{2}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{j}\|\nabla\Phi_{2}\|_{0}\\ &\quad+\|\Gamma_{2}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{0}\|\nabla\Phi_{2}\|_{j})\\ \leq&C(j)\delta^{1/2}\lambda^{(2\tau-1)(j-1)},\,\,j=0,1,2.\end{split}$
So combining with (3.26), (3.27), (3.37) and (3.38), we have
$\|v^{\star}-v\|_{0}+\|w^{\star}-w\|_{0}\leq$ $\|v^{\star}-v^{\lozenge}\|_{0}+\|v^{\lozenge}-v\|_{0}+\|w^{\star}-w^{\lozenge}\|_{0}+\|w^{\lozenge}-w\|_{0}$
$\leq$ $C(\delta^{1/2}\lambda^{-\tau}+\delta^{1/2}\lambda^{1-2\tau})\leq C\delta^{1/2}\lambda^{-\tau},$
$\|v^{\star}-v\|_{1}+\|w^{\star}-w\|_{1}\leq$ $C\delta^{1/2},$
which imply (3.11) and (3.12) respectively. Besides, from
$\|v^{\star}-v\|_{2}+\|w^{\star}-w\|_{2}\leq$ $\|v^{\star}-v^{\lozenge}\|_{2}+\|v^{\lozenge}-v\|_{2}+\|w^{\star}-w^{\lozenge}\|_{2}+\|w^{\lozenge}-w\|_{2}$
$\leq$ $C(\delta^{1/2}\lambda^{\tau}+\lambda^{2\tau-1})\leq C\delta^{1/2}\lambda^{2\tau-1},$
and (3.6) we also gets (3.13) since $\tau>1$.
Finally, we shall control the second deficit. A direct calculation implies
$\mathcal{E}_{2}:$ $=D(v^{\star},w^{\star})-(D(v^{\lozenge},w^{\lozenge})+a^{2}\nabla\Phi_{2}\otimes\nabla\Phi_{2})=\sum_{k=1}^{5}\mathcal{E}_{2k}$
with
$\mathcal{E}_{21}:=$ $\lambda^{1-2\tau}(\partial_{t}\Gamma_{1}\partial_{s}\Gamma_{1}\textrm{sym}(\nabla\Phi_{2}\otimes\nabla a)-\Gamma_{1}\nabla^{2}\tilde{v}^{\lozenge}+\partial_{s}\Gamma_{2}\textrm{sym}(\nabla\Phi_{1}\otimes\nabla a)),$
$\mathcal{E}_{22}:=$ ${\lambda}^{-\tau}\partial_{s}\Gamma_{1}\textrm{sym}((\nabla v^{\lozenge}-\nabla\tilde{v}^{\lozenge})\otimes\nabla a)+\partial_{t}\Gamma_{1}\textrm{sym}((\nabla v^{\lozenge}-\nabla\tilde{v}^{\lozenge})\otimes\nabla\Phi_{1}),$
$\mathcal{E}_{23}:=$ $\tfrac{1}{2}\lambda^{-2\tau}(\partial_{s}\Gamma_{1})^{2}\nabla a\otimes\nabla a,$
$\mathcal{E}_{24}:=$ $(V^{\star}-V^{\lozenge})\cdot\mathrm{Id},\,\,(\text{ here }\Delta V^{\star}=F(v^{\star}))$
$\mathcal{E}_{25}:=$ $((v^{\star}-v^{\lozenge})\cot\Theta)\cdot\mathrm{Id}.$
Repeating the estimating procedure in Step 2, we are able to get
(3.39) $\|\mathcal{E}_{2}\|_{0}\leq C\delta\lambda^{\tau-(2\tau-1)}=C\delta\lambda^{1-\tau},\quad\|\mathcal{E}_{2}\|_{1}\leq C\delta\lambda^{\tau}.$
In fact, when bounding the $C^{1}$ norm of the second matrix error, we only need to keep in mind that $\lambda^{2\tau-1}$(and $\lambda^{\tau}$) plays the same role as $\lambda^{\tau}$(and $\lambda$ repsectively) in Step 2. Therefore, the final matrix error can be calculated as follows
$\mathcal{E}$ $=D(v^{\star},w^{\star})-(D(v,w)+\rho^{2}(\textrm{Id}+H))$
$=D(v^{\star},w^{\star})-(D(v^{\lozenge},w^{\lozenge})+a^{2}\nabla\Phi_{2}\otimes\nabla\Phi_{2})$
$\qquad+D(v^{\lozenge},w^{\lozenge})-(D(v,w)+a^{2}\nabla\Phi_{1}\otimes\nabla\Phi_{1})+\tilde{h}-h$
$=\mathcal{E}_{1}+\mathcal{E}_{2}+\tilde{h}-h.$
After collecting the estimates of $\tilde{h}-h$ in Step 1, (3.35) and (3.39), we can achieve (3.14) and complete the proof.
4. Proof of Theorem 1.1 (2)
As discussed in the beginning of Section 3, we show Theorem 1.1 (2) by seeking infinitely many symmetric $2\times 2$ matrixes $A$’s and $C^{1,\alpha}(\alpha<\frac{1}{5})$ functions $v$’s such that
$\begin{cases}-\mbox{curl curl }A=1,\\ A=D(v,w)=\mathrm{sym}\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id},\end{cases}$
where $V$ solves the following Dirichlet problem:
(4.1) $\Delta V&=F(v)=2\nabla v\cdot\nabla\cot\Theta+v\Delta\cot\Theta,\mbox{ in }\Omega,\\ V&=0,\mbox{ on }\partial\Omega,\end{aligned}\right.$
We will solve the first system about $A$ together with a subsolution by seeking solutions to some Possion equations and the second system by constructing a sequence of adapted subsolutions via convex integration.
Proof of Theorem 1.1 (2).
We separate the proof into four steps.
Step 1. Matrix and initial subsolutions. We construct matrix $A$ and initial subsolutions at the same time. Let $u$ be the solution to the following Dirichlet problem,
$\Delta u&=0,\mbox{ in }\Omega,\\ u&=g,\mbox{ on }\partial\Omega,\end{aligned}\right.$
and $\psi$ be the solution of the following Dirichlet problem,
$-\Delta\psi&=1-\det\nabla^{2}u,\mbox{ in }\Omega,\\ \psi&=0,\mbox{ on }\partial\Omega.\end{aligned}\right.$
Define
$A=\psi\mathrm{Id}+\frac{1}{2}\nabla u\otimes\nabla u+(-u\cot\Theta+U)\mathrm{Id},$
where $U$ is the solution to
$\Delta U&=F(u):=2\nabla u\cdot\nabla\cot\Theta+u\Delta\cot\Theta,\mbox{ in }\Omega,\\ U&=0,\mbox{ on }\partial\Omega.\end{aligned}\right.$
Note that $|\sin\Theta|\geq c_{2}>0$ as in the assumptions of Theorem 1.1. So $\cot\Theta$ is a $C^{2,\kappa}$ bounded function. The existence and regularity of $u\in C^{2,\kappa}$, $\psi\in C^{2,\kappa}$ and $U\in C^{2,\kappa}$ are guaranteed by the elliptic theory.
Obviously, $A\in C^{1,\kappa}(\overline{\Omega})$. Besides,
$-\mathrm{curl}\,\mathrm{curl}\,A$ $=-\Delta\psi+\det\nabla^{2}u-\Delta(-u\cot\Theta+U)$
$=1+\Delta(u\cot\Theta)-F(u)$
$=1+\cot\Theta\Delta u=1.$
With $\Delta u=0$, we know $01-\det\nabla^{2}u=1+(\partial^{2}_{11}u)^{2}+(\partial^{2}_{12}u)^{2}>$ 0, and hence by the strong maximum principle, we have
$\psi(x)>0,\mbox{ for any }x\in\Omega.$
Set the initial subsolution
$(v_{0},w_{0})=(u,0).$
Then,
$A-D(v_{0},w_{0})=A-D(u,0)=\psi\mathrm{Id}:=\rho_{0}^{2}(\mathrm{Id}+H_{0}),$
where $\rho_{0}=\sqrt{\psi}$ and $H_{0}$ is the zero matrix. Moreover, $v_{0}|_{\partial\Omega}=u|_{\partial\Omega}=g$.
Note that $(v_{0},w_{0})$ and $A$ are not unique. For example, taking $v_{0}$ as before and $w_{0}\in\operatorname{\mathbb{R}}^{2}$ as arbitrary constant vector will also work.
Step 2. Iteration. With Proposition 3.1, for any $q\in\operatorname{\mathbb{N}}$, we construct inductively $(v_{q},w_{q},\rho_{q},H_{q})$ such that
  • $(1)_{q}$
    For any $x\in\overline{\Omega}$, it holds that $D_{q}=:A-D(v_{q},w_{q})=\rho_{q}^{2}(\textrm{Id}+H_{q}).$
  • $(2)_{q}$
    If $x\in\overline{\Omega}\setminus\tilde{\Omega}_{q}$, then $(v_{q},w_{q},\rho_{q},H_{q})(x)=(v,w,\rho,H)(x).$
  • $(3)_{q}$
    The following estimates hold in $\overline{\Omega}$:
    $\|\rho_{q}\|_{0}\leq 2\delta_{q+1}^{1/2},\quad\|H_{q}\|_{0}\leq 4\lambda_{q+1}^{-\frac{\alpha}{b}}.$
  • $(4)_{q}$
    If $x\in\tilde{\Omega}_{j}$ for some $j\geq q,$ we get the following estimates
    $\rho_{q}\geq\frac{3}{2}\delta_{j+2}^{1/2},\quad|\nabla\rho_{q}|\leq\delta_{j+1}^{1/2}\lambda_{j+2},\quad\frac{|\nabla\rho_{q}|}{\rho_{q}}\leq\lambda_{j+2},$
    $|\nabla^{2}v_{q}|+|\nabla^{2}w_{q}|\leq\delta_{j+1}^{1/2}\lambda_{j+2},\quad|\nabla H_{q}|\leq\lambda_{j+2}^{1-\frac{\alpha}{b}}.$
  • $(5)_{q}$
    For any $x\in\overline{\Omega},$ $q\geq 1,$
    (4.2) $\|v_{q}-v_{q-1}\|_{j}+\|w_{q}-w_{q-1}\|_{j}\leq\overline{C}\delta_{q}^{1/2}\lambda_{q}^{j-1},j=0,1,$
    where $\overline{C}$ is the constant in the conclusion of Proposition 3.1.
Here the amplitude sequence $\delta_{q}$ and the frequency sequence $\lambda_{q}$ are inductively defined by
$\delta_{1}=\max_{x\in\overline{\Omega}}\rho^{2},\quad\lambda_{q+1}=\lambda_{q}^{b},\quad\lambda_{q}=M\delta_{q}^{-\frac{1}{2\beta}},$
where arbitrary constant $\beta\in(0,\frac{1}{5}),$
(4.3) $b=1+\frac{9\sigma}{1-5\beta}\quad(\text{then }5\beta(b-1)+9\sigma=b-1).$
and $M$ is a large constant to be fixed.
Indeed, $(v_{0},w_{0},\rho_{0},H_{0})$ has been constructed in Step 1 and it is not hard to verify $(1)_{0}-(4)_{0}$. For $q\geq 1,$ using Proposition 3.1 and following Step 2 in the proof of Proposition 2.1 of [4], we are able to inductively get $(v_{q},w_{q},\rho_{q},H_{q}).$ Since the proof is exactly same, we omit the details.
Step 3. Taking limit. By $(3)_{q}$ and $(5)_{q}$, we know $(\rho_{q},H_{q})$ is a Cauchy sequence in $C^{0}(\overline{\Omega})$, and $(v_{q},w_{q})$ is a Cauchy sequence in $C^{1}(\overline{\Omega})$, respectively. Hence we can assume
$(v_{q},w_{q},\rho_{q},H_{q})\rightarrow(\bar{v},\bar{w},0,0)\text{ uniformly in }\overline{\Omega},$
for some functions $\bar{v}$ and $\bar{w}$.
Moreover, for any $\alpha\in(0,\frac{1}{5})$ and $\varepsilon>0$, we can choose $\beta\in(\alpha,\frac{1}{5})$ in Step 2, by $(4)_{q}-(5)_{q}$, $(4)_{q+1}$ and interpolation, we have for any $\alpha\in(0,\frac{\beta}{b^{2}})$,
$\|v_{q}-v_{q+1}\|_{1+\alpha}+\|w_{q}-w_{q+1}\|_{1+\alpha}\leq 4\delta_{q+1}^{(1-\alpha)/2}\delta_{q+2}^{\alpha/2}\lambda_{q+3}^{\alpha}\leq M^{\beta}\lambda_{q+1}^{\alpha b^{2}-\beta},$
which implies $(v_{q},w_{q})$ uniformly converges to $(\bar{v},\bar{w})$ in $C^{1,\alpha}(\overline{\Omega})$. Since $\alpha$ and $b$ can be taken arbitrarily close to 00 and $1$ by (4.3), respectively, $(\bar{v},\bar{w})\in C^{1,\alpha}({\Omega}),$ for any $\alpha<\beta$. From $(5)_{q}$, for any $\epsilon>0,$ we also have
$\|\bar{v}-v_{0}\|_{0}+\|\bar{w}-w_{0}\|_{0}\leq$ $\sum_{q=0}^{\infty}\|v_{q}-v_{q+1}\|_{0}+\|w_{q+1}-w_{q}\|_{0}$
$\leq$ $\sum_{q=0}^{\infty}\overline{C}\delta_{q+1}^{1/2}\lambda_{q+1}^{-1}\leq\overline{C}M^{-1}\leq\epsilon,$
if $M$ is taken larger.
Step 4. Very weak solution. We will show $\bar{v}$ is a very weak solution to (1.1). On the one hand, by construction, it is easy to see $(\bar{v},\bar{w})|_{\partial\Omega}=(v_{0},w_{0})|_{\partial\Omega}$. On the other hand, $\rho_{q}\to 0$ implies $A-D(\bar{v},\bar{w})=0.$ Thus for any $\phi\in C^{\infty}_{0}(\Omega)$, by integration by parts and curl crul $\mbox{curl crul }A=-1$, we have
$-\int_{\Omega}\phi\sin\Theta\mathrm{d}x=$ $\int_{\Omega}\mathrm{Trace}(H_{\phi\sin\Theta}\cdot A)\mathrm{d}x$
$=$ $\int_{\Omega}\mathrm{Trace}(H_{\phi\sin\Theta}\cdot\frac{1}{2}\nabla\bar{v}\otimes\nabla\bar{v})\mathrm{d}x-\int_{\Omega}\bar{v}\Delta(\phi\cos\Theta)\mathrm{d}x,$
where we have used curl $\mbox{curl curl}(\mathrm{sym}\nabla\bar{w})=0$ and notation
$H_{f}=\left(\begin{array}[]{rr}\partial_{22}^{2}f&-\partial_{12}^{2}f\\ -\partial_{12}^{2}f&\partial_{11}^{2}f\\ \end{array}\right)\text{ for any $C^{2}$ function }f.$
Then it implies
$0=\int_{\Omega}\partial_{1}\bar{v}\partial_{2}\bar{v}\partial_{12}^{2}(\phi\sin\Theta)$ $-\frac{1}{2}(\partial_{1}\bar{v})^{2}\partial_{22}^{2}(\phi\sin\Theta)-\frac{1}{2}(\partial_{2}\bar{v})^{2}\partial_{11}^{2}(\phi\sin\Theta)\mathrm{d}x$
$+\int_{\Omega}\bar{v}\Delta(\phi\cos\Theta)\mathrm{d}x-\int_{\Omega}\phi\sin\Theta\mathrm{d}x.$
Finally, since $\epsilon$ can be arbitrary and the initial subsolution is also not unique (note each $(v_{q},w_{q})$ can be an initial subsolution), (1.1) admits infinitely many very weak solutions $\bar{v}$’s. We then complete the proof. ∎
Appendix A Useful lemmas and proofs of Lemmas 3.2-3.3
In this appendix, we provide the sketch proofs of Lemmas 3.2-3.3 for completeness, which are as same as those in the first author [4]. We begin with the following two useful lemmas.
Lemma A.1 ([9]).
For any $r,s\geq 0$ and $0<\alpha\leq 1$, we have
$[h*\phi_{l}]_{r+s}\leq Cl^{-s}[h]_{r};$
$\|h-h*\phi_{l}\|_{r}\leq Cl^{1-r}[h]_{1},\mbox{ for }0\leq r\leq 1;$
$\|h-h*\phi_{l}\|_{j}\leq Cl^{2-j}\|h\|_{2},\mbox{ for }j=0,1;$
$\|(h_{1}h_{2})*\phi_{l}-(h_{1}*\phi_{l})(h_{2}*\phi_{l})\|_{r}\leq Cl^{2\alpha-r}\|h_{1}\|_{\alpha}\|h_{2}\|_{\alpha},$
with constant $C$ depending only on $s,r,\alpha,\phi$.
Lemma A.2 ([4]).
Let $\Sigma\subset\operatorname{\mathbb{R}}^{2}$ be a simply connected open bounded set with smooth boundary and $H:\Sigma\to\operatorname{\mathbb{R}}^{2\times 2}$ a smooth $2\times 2$ symmetric positive definite matrix valued function, such that for some $10<\alpha<$ 1, and $c,M\geq 1$,
$c^{-1}\mathrm{Id}\leq H\leq c\mathrm{Id},\,\,\|H\|_{C^{\alpha}(\Sigma)}\leq M.$
Then there exists a smooth diffeomorphism $\Phi:\Sigma\to\operatorname{\mathbb{R}}^{2}$ and a smooth positive function $\vartheta:\Sigma\to\operatorname{\mathbb{R}}$ satisfying
$H=\vartheta^{2}(\nabla\Phi_{1}\otimes\nabla\Phi_{1}+\nabla\Phi_{2}\otimes\nabla\Phi_{2}).$
Moreover the following estimates hold:
$\det(\nabla\Phi(x))\geq c_{0},\,\,\vartheta(x)\geq c_{0},\mbox{ for all }x\in\Sigma,$
$\|\vartheta\|_{j+\alpha}+\|\nabla\Phi\|_{j+\alpha}\leq C(j)\|H\|_{j+\alpha},\,\,j\in\operatorname{\mathbb{N}},$
where the constants $c_{0}>0$, $C(j)\geq 1$ depend only on $\alpha,c,M$ and on $\Sigma$.
Sketch proof of Lemma 3.2.
Recall the definition of $\Gamma_{i}:\operatorname{\mathbb{R}}\times\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}}$, $i=1,2$,
$\Gamma_{1}(s,t)=\frac{s}{\pi}\sin(2\pi t),\quad\Gamma_{2}(s,t)=-\frac{s^{2}}{4\pi}\sin(4\pi t).$
For any nonnegative integer $k$, it is not hard to show the following estimates.
(A.1) $\begin{split}|\partial_{t}^{k}\Gamma_{1}(s,t)|+|\partial_{s}\partial^{k}_{t}\Gamma_{2}(s,t)|&\leq C|s|,\\ |\partial_{s}\partial_{t}^{k}\Gamma_{1}(s,t)|&\leq C,\\ |\partial_{t}^{k}\Gamma_{2}(s,t)|&\leq Cs^{2}.\end{split}$
The derivatives of $\Gamma_{i}$, $i=1,2$ could be calculated directly by the chain rule, then the desired estimates are directly following by the assumptions on $\rho,H$, Lemma A.2 and (A.1). ∎
Sketch proof of Lemma 3.3.
By (3.15), (3.16), (3.17), (3.18), (3.24) and (3.25), we get
$\|\mathcal{E}_{11}\|_{0}\leq$ $\lambda^{-\tau}(\|\partial_{t}\Gamma_{1}\|_{0}\|\partial_{s}\Gamma_{1}\|_{0}\|\nabla a\|_{0}\|\nabla\Phi_{1}\|_{0}+\|\Gamma_{1}\|_{0}\|\nabla^{2}\tilde{v}\|_{0}$
$\quad+\|\partial_{s}\Gamma_{2}\|_{0}\|\nabla a\|_{0}\|\nabla\Phi_{1}\|_{0})\leq C\delta\lambda^{1-\tau},$
and
$\|\mathcal{E}_{12}\|_{0}\leq$ $\lambda^{-\tau}\|\partial_{s}\Gamma_{1}\|_{0}\|\nabla v-\nabla\tilde{v}\|_{0}\|\nabla a\|_{0}+\|\partial_{t}\Gamma_{1}\|_{0}\|\nabla v-\nabla\tilde{v}\|_{0}\|\nabla\Phi_{1}\|_{0}$
$\leq$ $C(\delta\lambda^{2-2\tau}+\delta\lambda^{1-\tau})\leq C\delta\lambda^{1-\tau},$
and
$\|\mathcal{E}_{13}\|_{0}\leq$ $\tfrac{1}{2}\lambda^{-2\tau}\|\partial_{s}\Gamma_{1}\|_{0}^{2}\|\nabla a\|_{0}^{2}\leq C\delta\lambda^{2-2\tau}.$
The gradients of $\operatorname{\mathcal{E}}_{11}$, $\operatorname{\mathcal{E}}_{12}$ and $\operatorname{\mathcal{E}}_{13}$ could be calculated directly by the chain rule, then the corresponding estimates can be got similarly. ∎

### v2

@@
-Ill-posedness of the Dirichlet problem for 2D Lagrangian mean curvature equation
+Ill-posedness of the Dirichlet problem for 2D Lagrangian mean curvature equations
 Abstract.
-We investigate the Dirichlet problem of the two dimensional Lagrangian mean curvature equation in a bounded domain.
+We investigate the Dirichlet problem of the two-dimensional Lagrangian mean curvature equation in bounded domains.
 Infinitely many $C^{1,\alpha}(\alpha\in(0,\frac{1}{5}))$ very weak solutions are built through Nash-Kuiper construction.
 Moreover, we note there are infinitely many $C^{1,\alpha}$ very weak solutions that can not be improved to be $C^{2,\alpha}$.
+2020 Mathematics Subject Classification:
+35A02, 35D99, 35J25, 35J60, 35R25
 1. Introduction
-In the present paper, we study the Dirichlet problem for two dimensional Lagrangian mean curvature equation
+In the present paper, we study the Dirichlet problem for the two-dimensional Lagrangian mean curvature equation
 (1.1) $\begin{cases}\cos\Theta\Delta v+\sin\Theta\left(\det\nabla^{2}v-1\right)=0\text{ in }\Omega,\\ v=g\text{ on }\partial\Omega,\end{cases}$
-where $\Omega\subset\operatorname{\mathbb{R}}^{2}$ is a bounded domain, phase function $\Theta:\overline{\Omega}\to(-\pi,\pi)$ is a given $C^{2,\kappa}(\kappa\in(0,1))$ function. The equation
+where $\Omega\subset\operatorname{\mathbb{R}}^{2}$ is a bounded domain, and the phase function $\Theta:\overline{\Omega}\to(-\pi,\pi)$ is a given $C^{2,\kappa}(\kappa\in(0,1))$ function. The equation
 (1.2) $\cos\Theta\Delta v+\sin\Theta\left(\det\nabla^{2}v-1\right)=0$
 has an equivalent form
-(1.3) $\arctan\lambda_{1}+\arctan\lambda_{2}=\Theta,$
+(1.3) $\arctan\lambda_{1}+\arctan\lambda_{2}=\Theta$
 with $\lambda_{1},\lambda_{2}$ being eigenvalues of $\nabla^{2}v$. The phase function $\Theta$ is a potential for the mean curvature of the Lagrangian submanifold
@@
 or called the gradient graph.
-Generally, $n$-dimensional Lagrangian mean curvature equation is
+Generally, the $n$-dimensional Lagrangian mean curvature equation is
 (1.4) $\sum_{k=1}^{n}\arctan\lambda_{k}=\Theta,$
 where $\lambda_{i}$’s are eigenvalues of $\nabla^{2}v$.
-The mean curvature vector of $n$-dimensional gradient graph $\Sigma$ is equal to $J\nabla_{g}\Theta$, where $g=\mathrm{Id}+(\nabla^{2}v)^{2}$ is the induced metric on $\Sigma$, and $J(x,y)=(-y,x)$ is a rotation in $\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n}$.
-In particular, when $\Theta$ is a constant, (1.4) is called the special Lagrangian equation, which originates in the special Lagrangian geometry introduced by Harvey-Lawson [14].
+The mean curvature vector of the $n$-dimensional gradient graph $\Sigma$ is equal to $J\nabla_{g}\Theta$, where $g=\mathrm{Id}+(\nabla^{2}v)^{2}$ is the induced metric on $\Sigma$, and J(x,y)=(y,x)J(x,y)=(-y,x) is a rotation in $\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n}$.
+In particular, when $\Theta$ is a constant, (1.4) is called the special Lagrangian equation, which originates in the special Lagrangian geometry introduced by Harvey-Lawson [HL].
 Note $\Sigma$ is a minimal surface in $\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n}$ if and only if $\Theta$ is a constant.
@@
 In this paper, we consider not only classical solutions but also very weak solutions to (1.1).
-In fact, very weak solution to the Monge-Ampère equation was first introduced by Lewicka-Pakzad [18] via very weak Hessian defined by Iwaniec [16].
-Two dimensional very weak Hessian coincides with the Monge-Ampère measure for any convex solution ([11]).
-Moreover, Without boundary conditions, Lewicka-Pakzad [18], Székelyhidi and the first author [7], and Hirsch-Inauen and the first author [5] proved that there are infinitely many very weak solutions to the two dimensional Monge-Ampère equation of different regularity.
-After imposing boundary data, the first author in [4] studied very weak solutions to the Monge-Ampère equation.
-Recently, Li-Qiu [19] also constructed infinitely many very weak solutions of the Dirichlet problem for 2-Hessian equation.
-All the above approaches to construct very weak solutions rely on the Nash-Kuiper construction, which is originated from isometric embeddings (Nash-Kuiper theorem, see [22, 17, 9, 6, 10, 23]), and developed to be convex integration to prove $h$-principle (see [13]).
-In recent decades, the technique of convex integration become a powerful approach to show non-uniqueness of nonlinear partial differential equations.
+In fact, the very weak solution to the Monge-Ampère equation was first introduced by Lewicka-Pakzad [LePa] via very weak Hessian defined by Iwaniec [Iwan].
+The two-dimensional very weak Hessian coincides with the Monge-Ampère measure for any convex solution ([DGG]).
+Moreover, Without boundary conditions, Lewicka-Pakzad [LePa], Székelyhidi and the first author [CSze], and Hirsch-Inauen and the first author [CHI] proved that there are infinitely many very weak solutions to the two dimensional Monge-Ampère equation of different regularity.
+After imposing boundary data, the first author in [Cao] studied very weak solutions to the Monge-Ampère equation.
+Recently, Li-Qiu [LQ] also constructed infinitely many very weak solutions of the Dirichlet problem for 2-Hessian equations.
+All the above approaches to constructing very weak solutions rely on the Nash-Kuiper construction, which was originated from isometric embeddings (the Nash-Kuiper theorem; see [Nash54, Kui55, CDS12, CS19-a, DIS, Sze]), and developed to be convex integration to prove the $h$-principle (see [Gro]).
+In recent decades, the technique of convex integration becomes a powerful approach to showing non-uniqueness of nonlinear partial differential equations.
 1.2. Main result and strategies
-Along the line of very weak solutions to the Monge-Ampère equation, we also introduce the concept of very weak solution to the Dirichlet problem (1.1).
+Along the line of very weak solutions to the Monge-Ampère equation, we also introduce the concept of very weak solutions to the Dirichlet problem (1.1).
 Definition 1.1 (Very weak solution).
-$v\in C^{1}(\Omega)$ is a very weak solution to the Dirichlet problem (1.1) in $\Omega$, if $v=g$ on $\partial\Omega$, and it solves the following equation in the very weak sense in $\Omega$,
-(1.5) $\cos\Theta\mbox{ curl curl }(v\cdot\mbox{Id})+\sin\Theta\left(-\frac{1}{2}\mbox{ curl curl }(\nabla v\otimes\nabla v)-1\right)=0,$
-namely, for any $\phi\in C^{\infty}_{0}(\Omega)$
+$v\in C^{1}(\Omega)$ is a very weak solution to the Dirichlet problem (1.1) in $\Omega$, if $v=g$ on $\partial\Omega$, and it satisfies the following equation in the very weak sense in $\Omega$:
+(1.5) $\cos\Theta\mbox{ curl curl }(v\cdot\mbox{Id})+\sin\Theta\left(-\frac{1}{2}\mbox{ curl curl }(\nabla v\otimes\nabla v)-1\right)=0;$
+namely, for any $\phi\in C^{\infty}_{0}(\Omega)$,
 $\int_{\Omega}\partial_{1}v\partial_{2}v\partial_{12}^{2}(\phi\sin\Theta)-\frac{1}{2}(\partial_{1}v)^{2}\partial_{22}^{2}(\phi\sin\Theta)-\frac{1}{2}(\partial_{2}v)^{2}\partial_{11}^{2}(\phi\sin\Theta)dx$
-[…]
-[…]
+(1.6) $\quad+\int_{\Omega}v\Delta(\phi\cos\Theta)dx-\int_{\Omega}\phi\sin\Theta dx=0.$
+The operator $\mathrm{curl}$ in two-dimension is defined by
+$\mbox{ curl }\xi$ $=\partial_{2}\xi_{1}-\partial_{1}\xi_{2}\mbox{ for }\xi=(\xi_{1},\xi_{2})\in\mathbb{R}^{2},$
+$\mbox{ curl }A$ $=(\partial_{2}A_{11}-\partial_{1}A_{12},\partial_{2}A_{21}-\partial_{1}A_{22})\mbox{ for }A=(A_{ij})_{2\times 2}\in\operatorname{\mathbb{R}}^{2\times 2}.$
+In particular,  $\mbox{ curl curl }(v\cdot\mbox{Id})$ is understood as $\Delta v$ in the very weak sense.
 Now we are ready to state our main result.
 Theorem 1.1.
-Let $\Omega\subset\operatorname{\mathbb{R}}^{2}$ be a bounded simply connected domain with $C^{2,\kappa}$ boundary $\partial\Omega$ for any given $\kappa\in(0,1)$.
+Let $\Omega\subset\operatorname{\mathbb{R}}^{2}$ be a bounded simply connected domain with a $C^{2,\kappa}$ boundary $\partial\Omega$ for any given $\kappa\in(0,1)$.
 Suppose $g\in C^{2,\kappa}(\partial\Omega)$ and the phase $\Theta:\overline{\Omega}\to(-\pi,\pi)$ is a $C^{2,\kappa}$ function.
@@
   • (1)
-    If $|\sin\Theta|\leq c_{1}$ for some small positive constant $c_{1}$, then the Dirichlet problem (1.1) admits a unique $C^{2,\kappa}(\Omega)$ classical solution.
+    There exists some small positive number $c_{1}$ such that if $|\sin\Theta|\leq c_{1}$, then the Dirichlet problem (1.1) admits a unique $C^{2,\kappa}(\Omega)$ classical solution.
   • (2)
-    If $|\sin\Theta|\geq c_{2}$ for any positive constant $c_{2}\in(0,1]$, then the Dirichlet problem (1.1) admits infinitely many $C^{1,\alpha}(\Omega)$ very weak solutions for any $\alpha\in(0,\frac{1}{5})$.
+    Given any constant $c_{2}\in(0,1]$, the Dirichlet problem (1.1) admits infinitely many $C^{1,\alpha}(\Omega)$ very weak solutions for any $\alpha\in(0,\frac{1}{5})$ whenever $|\sin\Theta|\geq c_{2}$.
 Remark 1.1.
-The assumption of simply connected is just from the application of Lemma A.2 in the proof of Proposition 3.1.
-Note $c_{2}$ is arbitrary, as a quick observation, we have
+The assumption of simply connectedness is only needed for applying Lemma A.2 to the proof of Proposition 3.1.
+Noting $c_{2}$ is arbitrary, as a quick observation, we have
 Corollary 1.2.
 For any $c_{3}\in(0,c_{1})$, if $c_{3}\leq|\sin\Theta|\leq c_{1}$, then there are infinitely many $C^{1,\alpha}$ very weak solutions to the Dirichlet problem (1.1) which can not be improved to be $C^{2,\alpha}$.
-Our strategy to show Theorem 1.1 (1) is considering (1.2) as a perturbation of Laplacian equation and then applying the implicit function theorem.
+Our strategy to show Theorem 1.1 (1) is considering (1.2) as a perturbation of a Laplacian equation and then applying the implicit function theorem.
 As for Theorem 1.1 (2), with $|\sin\Theta|\geq c_{2}>0$, the problem is reduced to solve
@@
 We will solve the above equation by the Nash-Kuiper construction.
-The main obstacle is how to treat terms with phase function, comparing with [4], which will be regarded as a Laplacian of some function, namely,
-$\cot\Theta\Delta v=\Delta(v\cot\Theta-V),\text{ in the very weak sense,}$,
+The main obstacle is how to treat terms with the phase function, compared with [Cao], which will be regarded as a Laplacian of some function, i.e.,
+$\cot\Theta\Delta v=\Delta(v\cot\Theta-V),\text{ in the very weak sense,}$
 with
@@
 Hence, the problem is turned to solve
-curlcurl( sym $\mathrm{curl}\,\mathrm{curl}\,\big{(}\mbox{ sym }\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id}\big{)}=-1,$
-in the very weak sense, and then $V\cdot\textrm{Id}$ becomes an error term of lower order. With the elliptic estimate in Lemma 3.4, we are able to control it and arrive at our conclusion.
-We remark that the method proving Theorem 1.1 (2) has more general application.
+$\mathrm{curl}\,\mathrm{curl}\,\big(\mbox{ sym }\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id}\big)=-1,$
+in the very weak sense, and then $V\cdot\textrm{Id}$ becomes an error term of a lower order. With the elliptic estimate in Lemma 3.4, we are able to control it and arrive at our conclusion.
+We remark that the method proving Theorem 1.1 (2) has more general applications.
 Remark 1.2.
@@
 $F_{\tau}(\lambda)=\left\{\begin{aligned} &\frac{1}{2}\big(\ln\lambda_{1}+\ln\lambda_{2}\big),\qquad\tau=0,\\ &\frac{\sqrt{a^{2}+1}}{2b}\big(\ln\frac{\lambda_{1}+a-b}{\lambda_{1}+a+b}+\ln\frac{\lambda_{2}+a-b}{\lambda_{2}+a+b}\big),\qquad 0<\tau<\frac{\pi}{4},\\ &-\sqrt{2}\big(\frac{1}{1+\lambda_{1}}+\frac{1}{1+\lambda_{2}}\big),\qquad\tau=\frac{\pi}{4},\\ &\frac{\sqrt{a^{2}+1}}{b}\big(\arctan\frac{\lambda_{1}+a-b}{\lambda_{1}+a+b}+\arctan\frac{\lambda_{2}+a-b}{\lambda_{2}+a+b}\big),\qquad\frac{\pi}{4}<\tau<\frac{\pi}{2},\\ &\arctan\lambda_{1}+\arctan\lambda_{2},\qquad\tau=\frac{\pi}{2},\end{aligned}\right$.
-where $a=\cot\tau$, $b=\sqrt{|\cot^{2}\tau-1|}$. With slight modifications, we can establish result like Theorem 1.1 (2) to the following Dirichlet problem
+where $a=\cot\tau$, and $b=\sqrt{|\cot^{2}\tau-1|}$. With slight modifications, we can establish the result like Theorem 1.1 (2) for the following Dirichlet problem
 (1.7) $\begin{cases}F_{\tau}(\lambda)=\Theta,&\mbox{ in }\Omega,\\ \qquad v=g,&\mbox{ on }\partial\Omega,\end{cases}$
-where $\lambda_{1}$, $\lambda_{2}$ are eigenvalues of $\nabla^{2}v$. We note the equations $F_{\tau}(\lambda)=\Theta$ can be formulated to equations of the following form
+where $\lambda_{1}$, $\lambda_{2}$ are eigenvalues of $\nabla^{2}v$. We note the equation $F_{\tau}(\lambda)=\Theta$ can be formulated as the equation of the following form
 $P\Delta u+Q\det\nabla^{2}u=R.$
-Equations in (1.7) are originated from the study of gradient graphs $\Sigma$.
-By Warren [26], if $\Sigma$ is a minimal Lagrangian graph in $(\operatorname{\mathbb{R}}^{2}\times\operatorname{\mathbb{R}}^{2},g_{\tau})$ with
-$g_{\tau}=\sin\tau\big{(}\sum_{i=1}^{2}\mathrm{d}x_{i}\otimes\mathrm{d}x_{i}+\sum_{j=1}^{2}\mathrm{d}y_{j}\otimes\mathrm{d}y_{j}\big{)}+\cos\tau\big{(}\sum_{i=1}^{2}\mathrm{d}x_{i}\otimes\mathrm{d}y_{i}+\sum_{j=1}^{2}\mathrm{d}y_{j}\otimes\mathrm{d}x_{j}\big{)},$
+Equations in (1.7) are originated from the study of the gradient graph $\Sigma$.
+By Warren [Wa], if $\Sigma$ is a minimal Lagrangian graph in $(\operatorname{\mathbb{R}}^{2}\times\operatorname{\mathbb{R}}^{2},g_{\tau})$ with
+$g_{\tau}=\sin\tau\big(\sum_{i=1}^{2}dx_{i}\otimes dx_{i}+\sum_{j=1}^{2}dy_{j}\otimes dy_{j}\big)+\cos\tau\big(\sum_{i=1}^{2}dx_{i}\otimes dy_{i}+\sum_{j=1}^{2}dy_{j}\otimes dx_{j}\big),$
 then
-$F_{\tau}(\lambda(\nabla^{2}v))=\text{constant},$
-We refer readers to results about gradient graphs in general dimensions in [26].
+$F_{\tau}(\lambda(\nabla^{2}v))=\text{constant}.$
+We refer readers to the results about gradient graphs in general dimensions in [Wa].
 1.3. Organization and notations
@@
 where the Hölder semi-norm is defined by
-$[h]_{m+\alpha}=\max_{|\beta|=m}\sup_{x\neq y}\frac{|\partial^{\beta}h(x)-\partial^{\beta}h(y)|}{|x-y|^{\alpha}}.$
+$h]_{m+\alpha}=\max_{|\beta|=m}\sup_{x\neq y}\frac{|\partial^{\beta}h(x)-\partial^{\beta}h(y)|}{|x-y|^{\alpha}}.$
 Here $m$ is a nonnegative integer, $\alpha\in(0,1]$ a constant, and $\beta$ be a multi-index.
-We denote by $B_{r}$ the ball in $\operatorname{\mathbb{R}}^{2}$ with radius $r$ centering at 00.
-Let $\varphi\in C^{\infty}_{0}(B_{1})$ be a standard two dimensional mollifier.
+We denote by $B_{r}$ the ball in $\operatorname{\mathbb{R}}^{2}$ with radius $r$ centering at 0.
+Let $\varphi\in C^{\infty}_{0}(B_{1})$ be a standard two-dimensional mollifier.
 We define $\varphi_{l}(x)=l^{-2}\varphi(l^{-1}x)$, and $l>0$ is called the length-scale.
-We also denote
-[…]
-Some properties of mollification can be found in Lemma A.1. The operator $\mathrm{curl}$ in two dimension is defined by
-[…]
-[…]
-Besides, for $\xi=(\xi_{1},\xi_{2})\in\mathbb{R}^{2}$ and $\eta=(\eta_{1},\eta_{2})\in\mathbb{R}^{2}$, we write $\xi\otimes\eta$ a $2\times 2$-matrix with entries $(\xi\otimes\eta)_{ij}=\xi_{i}\eta_{j}$.
+We also define
+$\tilde{h}(x)=h*\varphi_{l}(x)=l^{-2}\int_{\operatorname{\mathbb{R}}^{2}}h(x-y)\varphi(\frac{y}{l})dy.$
+Some properties of mollification can be found in Lemma A.1.
+Besides, for $\xi=(\xi_{1},\xi_{2})\in\mathbb{R}^{2}$ and $\eta=(\eta_{1},\eta_{2})\in\mathbb{R}^{2}$, we write $\xi\otimes\eta$ as a $2\times 2$-matrix with entries $(\xi\otimes\eta)_{ij}=\xi_{i}\eta_{j}$.
 By a direct calculation, for a $C^{3}$-function $w$, we have
@@
 $\mathrm{sym}\nabla w=\frac{1}{2}(\nabla w+\nabla w^{T}).$
-Acknowledgement
-Wentao Cao’s research is supported in part by the National Natural Science Foundation of China (Grant No.
+Acknowledgements
+Wentao Cao’s research was supported in part by National Natural Science Foundation of China (Grant No.
 12471224).
-Zhehui Wang’s research is supported in part by Young Scientists Fund of the National Natural Science Foundation of China (Grant No.
+Zhehui Wang’s research was supported in part by Young Scientists Fund of National Natural Science Foundation of China (Grant No.
 12401247), and Guangdong Basic and Applied Basic Research Foundation (Grant No.
@@
 2. Proof of Theorem 1.1 (1)
-The proof of Theorem 1.1 is based on an application of the implicit function theorem.
+The proof of Theorem 1.1 (1) is based on the application of the implicit function theorem.
 Proof.
@@
 Then we have the following conclusion.
-Claim: There is a unique $\psi\in\mathcal{X}$ and a small enough constant $\mu\in(0,1)$ such that $G(\psi,\Psi)=0$ for any $\Psi$ satisfying $\|\Psi\|_{\kappa}\leq\mu$.
+Claim: There are a unique $\psi\in\mathcal{X}$ and a small enough constant $\mu\in(0,1)$ such that $G(\psi,\Psi)=0$ for any $\Psi$ satisfying $\|\Psi\|_{\kappa}\leq\mu$.
 In fact, the Fréchet derivative of $G$ with respect to $\psi$ is given by
@@
 $\|f\|_{C^{\kappa}(B_{r})}\leq\sigma\|f\|_{C^{1,\kappa}(B_{r})}+(\frac{C}{\sigma^{\kappa}}+1)\|f\|_{L^{\infty}(B_{r})}$
-for any $0<\sigma\leq r$ and $f\in C^{1,\kappa}(B_{r})$. Hence, if $|\sin\Theta|\leq c_{1}$ with small constant $c_{1}$ to be fixed, then we can get
+for any $0<\sigma\leq r$ and $f\in C^{1,\kappa}(B_{r})$. Hence, if $|\sin\Theta|\leq c_{1}$ with the small constant $c_{1}$ to be fixed, then we can get
 $\|\sin\Theta\|_{\kappa}\leq C(\kappa,\|\Theta\|_{1+\kappa})(c_{1}+c_{1}^{1-\kappa}),$
@@
 $\Delta\psi+\tan\Theta(\det\nabla^{2}(\psi+\widetilde{g})-1)+\Delta\widetilde{g}=0.$
-Then, $v:=\psi+\widetilde{g}$ uniquely solves
-$&\Delta v+\tan\Theta(\det\nabla^{2}v-1)=0\mbox{ in }\Omega,\\ &v=g\mbox{ on }\partial\Omega.\end{aligned}\right.$
+Then, $v:=\psi+\widetilde{g}$ uniquely satisfies
+$\left\{\begin{aligned} &\Delta v+\tan\Theta(\det\nabla^{2}v-1)=0\mbox{ in }\Omega,\\ &v=g\mbox{ on }\partial\Omega.\end{aligned}\right$.
 Hence, we complete the proof of Theorem 1.1 (1). ∎
-3. Stage Proposition
-The rest of the paper is devoted to show Theorem 1.1 (2). Similar to the treatment in [4, 7, 18] for the Monge-Ampère equation, we split (1.2) (after dividing $\sin\Theta$) into a differential system
-(3.1) curl curl $-\mbox{curl curl }A=1,$
+3. The Stage Proposition
+The rest of the paper is devoted to showing Theorem 1.1 (2).
+Similar to the treatment in [Cao, CSze, LePa] for the Monge-Ampère equation, we split (1.2) (after dividing $\sin\Theta$) into a differential system
+(3.1) $-\mbox{curl curl }A=1,$
 (3.2) $A=\mbox{ sym }\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id},$
 where $V$ is the solution to
-$\Delta V&=F(v),\mbox{ in }\Omega,\\ V&=0,\mbox{ on }\partial\Omega,\end{aligned}\right.$
+$\left\{\begin{aligned} \Delta V&=F(v),\mbox{ in }\Omega,\\ V&=0,\mbox{ on }\partial\Omega,\end{aligned}\right$.
 and $F(v)$ is defined as
 (3.3) $F(v)=2\nabla v\cdot\nabla\cot\Theta+v\Delta\cot\Theta.$
-In fact, by a direct calculation, we note that the $v$ satisfying (3.1)-(3.2) solves (1.5) in the very weak sense.
-To solve (3.2), our strategy is applying the technique of Nash-Kuiper construction to construct a sequence of subsolutions satisfying boundary condition, whose limit is our desired solution. Set
+In fact, by a direct calculation, we note that $v$ satisfying (3.1)-(3.2) is a solution of the equation (1.5) in the very weak sense.
+To solve (3.2), our strategy is applying the technique of Nash-Kuiper construction to construct a sequence of subsolutions satisfying the boundary condition, whose limit is our desired solution. Set
 (3.4) $D(v,w):=\mbox{ sym }\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id},$
@@
 where $D(v,w)$ is as (3.4), and $\rho\in C(\overline{\Omega})$ is a nonnegative function with $\partial\Omega\subset\{x:\rho(x)=0\}$, and $H\in C(\overline{\Omega};\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ is a symmetric matrix such that $\textrm{Id}+H$ is positive definite.
-Note that the definition is a slight modification of that in [4].
-To construct new adapted subsolutions, we need to add the deficit $A-D(v,w)$ to the given adapted subsolution $v.$ Since the deficit is of the form $\rho^{2}(\mathrm{Id}+H)$, we require the following stage proposition (in the terminology of Nash [22]) for iteration, which is parallel to Proposition 3.1 in [4] and whose proof is different due to the new error terms.
+Note that the definition is a slight modification of that in [Cao].
+To construct new adapted subsolutions, we need to add the deficit AD(v,w)A-D(v,w) to the given adapted subsolution $v.$ Since the deficit is of the form $\rho^{2}(\mathrm{Id}+H)$, we require the following stage proposition (in the terminology of Nash [Nash54]) for iteration, which is parallel to Proposition 3.1 in [Cao] and whose proof is different due to the new error terms.
 The new error terms are then controlled through Lemma 3.4.
 Proposition 3.1.
-Suppose $\Sigma\subset\operatorname{\mathbb{R}}^{2}$ is a simply connected open bounded set with $C^{2}$ boundary.
+Suppose $\Sigma\subset\operatorname{\mathbb{R}}^{2}$ is a simply connected open bounded set with a $C^{2}$ boundary.
 Let $v\in C^{2}(\overline{\Sigma})$ and $w\in C^{2}(\Sigma,\operatorname{\mathbb{R}}^{2})$ such that
@@
 (3.6) $\|v\|_{2}+\|w\|_{2}\leq\delta^{\frac{1}{2}}\lambda,$
-and let $\rho\in C^{1}(\overline{\Sigma})$, $H\in C^{1}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ such that
+and let $\rho\in C^{1}(\overline{\Sigma})$, $H\in C^{1}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ be such that
 (3.7) $\|\rho\|_{0}\leq\delta^{\frac{1}{2}},\quad\|\nabla\rho\|_{0}\leq\delta^{\frac{1}{2}}\lambda,$
 (3.8) $\|H\|_{0}\leq\lambda^{-\alpha},\quad\|\nabla H\|_{0}\leq\lambda^{1-\alpha},$
-for some $\gamma>1$, $10<\delta,\,\alpha<$ 1 and $\lambda>1$ with $2\leq\lambda^{\alpha}$ and $\delta^{1/2}\lambda>1$.
-Then for any $\tau>1$, there exist functions $v^{\star}\in C^{2}(\overline{\Sigma})$, $w^{\star}\in C^{2}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2})$ and symmetric matrix $\operatorname{\mathcal{E}}\in C^{1}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ such that
+for some $\gamma>1$, $0<\delta,\,\alpha<1$ and $\lambda>1$ with $2\leq\lambda^{\alpha}$ and $\delta^{1/2}\lambda>1$.
+Then for any $\tau>1$, there exist functions $v^{\star}\in C^{2}(\overline{\Sigma})$, $w^{\star}\in C^{2}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2})$ and a symmetric matrix $\operatorname{\mathcal{E}}\in C^{1}(\overline{\Sigma},\operatorname{\mathbb{R}}^{2\times 2}_{sym})$ such that
 (3.9) $D(v^{\star},w^{\star})=D(v,w)+\rho^{2}(\mathrm{Id}+H)+\operatorname{\mathcal{E}},\,\,\mbox{in }\overline{\Sigma},$
@@
 where constant $C$ and $\bar{C}$ depend only on $\gamma$ and $\alpha$.
-Comparing the proof of Proposition 3.1 in [4], the key ingredient and difference of the proof for our Proposition 3.1 is applying the elliptic estimate in Lemma 3.4 to control the new error terms with the phase function.
+Our proof for the proposition above differs from that for Proposition 3.1 in [Cao] due to the presence of the phase function $\Theta$.
+Indeed, the key ingredients of our proof consist of applying the elliptic estimates in Lemma 3.4 to control the error terms arising from $\Theta$.
 Proof.
-Same as [4], we also divide the proof into three steps.
-Step 1. Mollification. Take the lengthscale $\ell=\lambda^{-\tau}$, Set $\tilde{\rho}=\rho*\varphi_{\ell}$ and $\tilde{H}=H*\varphi_{\ell}$. Then by (3.7), (3.8) and Lemma A.1, we know,
+We divide the proof into three steps.
+Step 1. Mollification. Take the lengthscale $\ell=\lambda^{-\tau}$. Set $\tilde{\rho}=\rho*\varphi_{\ell}$ and $\tilde{H}=H*\varphi_{\ell}$. Then by (3.7), (3.8) and Lemma A.1, we know,
 $\|\tilde{\rho}\|_{0}\leq C\delta^{\frac{1}{2}},\quad\|\tilde{\rho}\|_{j}\leq C(j)\ell^{1-j}[\rho]_{1}\leq C(j)\delta^{\frac{1}{2}}\lambda\ell^{1-j},$
@@
 $\|\tilde{H}-H\|_{0}\leq C\ell[H]_{1}\leq C\lambda^{1-\alpha}\ell,$
-for $j\in\operatorname{\mathbb{N}}_{+}$. Let $\tilde{h}=\tilde{\rho}^{2}(\mathrm{Id}+\tilde{H})$ and $h=\rho^{2}(\mathrm{Id}+{H})$, then
+[…]
 $\|\tilde{h}-h\|_{0}$ $\leq\|\tilde{\rho}^{2}-\rho^{2}\|_{0}+\|\tilde{\rho}^{2}\tilde{H}-\rho^{2}H\|_{0}\leq C\delta\lambda\ell+C\delta\lambda^{1-\alpha}\ell\leq C\delta\lambda^{1-\tau},$
@@
 Besides, it holds that
-$+\tilde{H}\geq(1-\lambda^{-\alpha})\textrm{Id}>\tfrac{1}{2}\textrm{Id},$ Id ,
+$\textrm{Id}+\tilde{H}\geq(1-\lambda^{-\alpha})\textrm{Id}>\tfrac{1}{2}\textrm{Id},$
 and
@@
 (3.17) $\|\tilde{v}-v\|_{1}+\|\tilde{w}-w\|_{1}\leq C\ell(\|v\|_{2}+\|w\|_{2})\leq C\delta^{\frac{1}{2}}\lambda^{1-\tau},$
-(3.18) $\|\tilde{v}\|_{2+j}+\|\tilde{w}\|_{2+j}\leq C\ell^{-j}(\|v\|_{2}+\|w\|_{2})\leq C\delta^{\frac{1}{2}}\lambda^{1+\tau j},$
+(3.18) $\|\tilde{v}\|_{2+j}+\|\tilde{w}\|_{2+j}\leq C\ell^{-j}(\|v\|_{2}+\|w\|_{2})\leq C\delta^{\frac{1}{2}}\lambda^{1+\tau j}.$
 Define
@@
 $+\lambda^{-\tau}\Gamma_{2}(a(x),\lambda^{\tau}\Phi_{1}(x))\nabla\Phi_{1}(x),$
-where $\Gamma_{i},i=1,2$ are defined as follows.
+where $\Gamma_{i},i=1,2$ are defined as follows:
 (3.19) $\Gamma_{1}(s,t)=\frac{s}{\pi}\sin(2\pi t),\quad\Gamma_{2}(s,t)=-\frac{s^{2}}{4\pi}\sin(4\pi t).$
@@
 (3.23) $\begin{split}\|w^{\lozenge}-w\|_{j}\leq C(j)\lambda^{-\tau}(&\|\Gamma_{1}\|_{j}\|\tilde{v}\|_{1}+\|\Gamma_{1}\|_{0}\|\tilde{v}\|_{j+1}\\ &+\|\Gamma_{2}\|_{j}\|\nabla\Phi_{1}\|_{0}+\|\Gamma_{2}\|_{0}\|\nabla\Phi_{1}\|_{j}),\end{split}$
-for $j=0,1,2.$ Here $\|\Gamma_{i}\|_{j}$ denotes $C^{j}$ norms of function $x\rightarrow\Gamma_{i}(a(x),\lambda^{\tau}\Phi_{1}(x)).$
+for $j=0,1,2.$ Here $\|\Gamma_{i}\|_{j}$ denotes the $C^{j}$ norms of the function $x\rightarrow\Gamma_{i}(a(x),\lambda^{\tau}\Phi_{1}(x)).$
 Lemma 3.2.
@@
 (3.25) $\begin{split}&\|\Gamma_{2}\|_{0}+\|\partial_{t}\Gamma_{2}\|+\|\partial_{t}^{2}\Gamma_{2}\|_{0}\leq C\delta,\\ &\|\Gamma_{2}\|_{1}+\|\partial_{t}\Gamma_{2}\|_{1}\leq C\delta\lambda^{\tau},\\ &\|\partial_{s}\Gamma_{2}\|_{j}\leq C\delta^{1/2}\lambda^{\tau j},\,j=0,1,\\ &\|\Gamma_{2}\|_{2}\leq C\delta\lambda^{2\tau}.\end{split}$
-The proof of Lemma 3.2 is same as that in [4], but for the completeness, we put it in Appendix.
+The proof of Lemma 3.2 is the same as that in [Cao], but for the completeness, we put it in Appendix.
 So from (3.22), (3.15), (3.23), and Lemma 3.2, we get
-(3.26) $\|v^{\lozenge}-v\|_{j}\leq C\delta^{1/2}\lambda^{\tau j-\tau},\,\,j=0,1,2.$
+(3.26) $\|v^{\lozenge}-v\|_{j}\leq C\delta^{1/2}\lambda^{\tau j-\tau},\,\,j=0,1,2,$
 and
@@
 (3.28) $\begin{split}&\|v^{\lozenge}\|_{1}+\|w^{\lozenge}\|_{1}\leq C,\\ &\|v^{\lozenge}\|_{2}+\|w^{\lozenge}\|_{2}\leq C\delta^{1/2}\lambda^{\tau}.\end{split}$
-Next we shall control $C^{j},j=0,1$ norms of the error matrix $\mathcal{E}.$ Taking derivatives gives
+Next we shall control the $C^{j}$($j=0,1$) norms of the error matrix $\mathcal{E}.$ Taking derivatives gives
 $\nabla v^{\lozenge}$ $=\nabla v+\lambda^{-\tau}\partial_{s}\Gamma_{1}\nabla a+\partial_{t}\Gamma_{1}\nabla\Phi_{1},$
@@
 Hence, with (3.20), we have
-$\mathcal{E}_{1}=$ $D(v^{\lozenge},w^{\lozenge})-(D(v,w)+a^{2}\nabla\Phi_{1}\otimes\nabla\Phi_{1})=\sum_{k=1}^{5}\mathcal{E}_{1k}$
+$\mathcal{E}_{1}=$ $D(v^{\lozenge},w^{\lozenge})-\left[D(v,w)+a^{2}\nabla\Phi_{1}\otimes\nabla\Phi_{1}\right]=\sum_{k=1}^{5}\mathcal{E}_{1k}$
 with
-$\mathcal{E}_{11}:=$ $\lambda^{-\tau}(\partial_{t}\Gamma_{1}\partial_{s}\Gamma_{1}\mathrm{sym}(\nabla\Phi_{1}\otimes\nabla a)-\Gamma_{1}\nabla^{2}\tilde{v}+\partial_{s}\Gamma_{2}\textrm{sym}(\nabla\Phi_{1}\otimes\nabla a)),$
-$\mathcal{E}_{12}:=$ ${\lambda}^{-\tau}\partial_{s}\Gamma_{1}\textrm{sym}((\nabla v-\nabla\tilde{v})\otimes\nabla a)+\partial_{t}\Gamma_{1}\textrm{sym}((\nabla v-\nabla\tilde{v})\otimes\nabla\Phi_{1}),$
+$\mathcal{E}_{11}:=$ $\lambda^{-\tau}\left[\partial_{t}\Gamma_{1}\partial_{s}\Gamma_{1}\mathrm{sym}(\nabla\Phi_{1}\otimes\nabla a)-\Gamma_{1}\nabla^{2}\tilde{v}+\partial_{s}\Gamma_{2}\textrm{sym}(\nabla\Phi_{1}\otimes\nabla a)\right],$
+$\mathcal{E}_{12}:=$ ${\lambda}^{-\tau}\partial_{s}\Gamma_{1}\textrm{sym}\left[(\nabla v-\nabla\tilde{v})\otimes\nabla a\right]+\partial_{t}\Gamma_{1}\textrm{sym}\left[(\nabla v-\nabla\tilde{v})\otimes\nabla\Phi_{1}\right],$
 $\mathcal{E}_{13}:=$ $\tfrac{1}{2}\lambda^{-2\tau}(\partial_{s}\Gamma_{1})^{2}\nabla a\otimes\nabla a,$
 $\mathcal{E}_{14}:=$ $(V^{\lozenge}-V)\cdot\mathrm{Id},\,\,(\text{ here }\Delta V^{\lozenge}=F(v^{\lozenge}))$
-$\mathcal{E}_{15}:=$ $((v^{\lozenge}-v)\cot\Theta)\cdot\mathrm{Id}$
-Estimates of the first three errors and their proof are same as those in [4].
+$\mathcal{E}_{15}:=$ $\left[(v^{\lozenge}-v)\cot\Theta\right]\cdot\mathrm{Id}$
+Estimates of the first three errors and their proofs are the same as those in [Cao].
 Lemma 3.3.
-Assume all the assumptions of Proposition 3.1 are fulfilled, then for $j=0,1$, the following estimates hold.
+Assume all the assumptions of Proposition 3.1 are fulfilled, then for $j=0,1$, the following estimates hold:
 $\|\mathcal{E}_{11}\|_{j}\leq C\delta\lambda^{1-\tau(1-j)},$
@@
 Note that
-$\Delta(V^{\lozenge}-V)=F(v^{\lozenge})-F(v)&=\operatorname{div}\zeta+f,\mbox{ in }\Omega,\\ V^{\lozenge}-V&=0,\mbox{ on }\partial\Omega.\end{aligned}\right.$
-Hence we require the follow elliptic estimates.
+$\left\{\begin{aligned} \Delta(V^{\lozenge}-V)=F(v^{\lozenge})-F(v)&=\operatorname{div}\zeta+f,\mbox{ in }\Omega,\\ V^{\lozenge}-V&=0,\mbox{ on }\partial\Omega.\end{aligned}\right$.
+Hence we require the following elliptic estimates.
 Lemma 3.4.
@@
 (3.31) $\|u_{2}\|_{0}\leq C(n,\mathrm{diam}(\Omega))\|f\|_{0},$
-where $\mathrm{diam}(\Omega)$ is the diameter of $\Omega$. Now (3.29) will follow by (3.30), (3.31) and the claim below:
+where $\mathrm{diam}(\Omega)$ is the diameter of $\Omega$. Now (3.29) will follow from (3.30), (3.31) and the claim below:
 Claim: $\|u_{1}\|_{0}\leq C(n,\Omega)\|\zeta\|_{0}.$
@@
 then
-$\int_{\Omega}u_{1}^{2}\mathrm{d}x$ $=\int_{\Omega}u_{1}\Delta\phi\mathrm{d}x=\int_{\Omega}\phi\Delta u_{1}\mathrm{d}x=\int_{\Omega}\phi\operatorname{div}\zeta\mathrm{d}x$
-$=-\int_{\Omega}\nabla\phi\cdot\zeta\mathrm{d}x\leq\|\nabla\phi\|_{L^{2}}\|\zeta\|_{L^{2}}.$
+$\int_{\Omega}u_{1}^{2}dx$ $=\int_{\Omega}u_{1}\Delta\phi dx=\int_{\Omega}\phi\Delta u_{1}dx=\int_{\Omega}\phi\operatorname{div}\zeta dx$
+$=-\int_{\Omega}\nabla\phi\cdot\zeta dx\leq\|\nabla\phi\|_{L^{2}}\|\zeta\|_{L^{2}}.$
 Besides, the maximum principle tells us that
 (3.32) $\|\phi\|_{0}\leq C(n,\mathrm{diam}(\Omega))\|u_{1}\|_{0}.$
-Note
-$\int_{\Omega}|\nabla\phi|^{2}\mathrm{d}x=-\int_{\Omega}\phi\Delta\phi\mathrm{d}x=-\int_{\Omega}\phi u_{1}\mathrm{d}x\leq\|\phi\|_{L^{2}}\|u_{1}\|_{L^{2}},$
+Noting
+$\int_{\Omega}|\nabla\phi|^{2}dx=-\int_{\Omega}\phi\Delta\phi dx=-\int_{\Omega}\phi u_{1}dx\leq\|\phi\|_{L^{2}}\|u_{1}\|_{L^{2}},$
 we have
@@
 (3.33) $\|u_{1}\|^{\frac{3}{2}}_{L^{2}}\leq\|\phi\|_{L^{2}}^{\frac{1}{2}}\|\zeta\|_{L^{2}}\leq|\Omega|^{\frac{1}{4}}\|\phi\|_{0}^{\frac{1}{2}}|\Omega|^{\frac{1}{2}}\|\zeta\|_{0}\leq C(n,\Omega)\|u_{1}\|^{\frac{1}{2}}_{0}\|\zeta\|_{0}.$
-here we apply (3.32) in last inequality. By Theorem 8.15 in [12], we know
+Here we apply (3.32) in the last inequality. By Theorem 8.15 in [GT], we know
 $\|u_{1}\|_{0}\leq C(n,\Omega)(\|\zeta\|_{0}+\|u_{1}\|_{L^{2}}),$
@@
 (3.34) $\|u_{1}\|_{0}\leq C(n,\Omega)(\|\zeta\|_{0}+\|u_{1}\|_{0}^{\frac{1}{3}}\|\zeta\|_{0}^{\frac{2}{3}}).$
-Denote $D=\|u_{1}\|_{0}$ and $B=\|\zeta\|_{0}$. We consider the following two cases.
+Let $D=\|u_{1}\|_{0}$ and $B=\|\zeta\|_{0}$. We consider the following two cases.
 Case 1: $D^{\frac{1}{3}}B^{\frac{2}{3}}\leq B$. In this case, we have $D\leq B$ and the claim follows.
@@
 $\|V^{\lozenge}-V\|_{0}\leq C(\|f\|_{0}+\|\zeta\|_{0})\leq C\|v^{\lozenge}-v\|_{0}\leq C\delta^{\frac{1}{2}}\lambda^{-\tau},$
-and for any $\sigma\in(0,1)$, by the $C^{1,\sigma}$ estimates ((4.45)-(4.46) in [12]) ,
+and for any $\sigma\in(0,1)$, by the $C^{1,\sigma}$ estimates ((4.45)-(4.46) in [GT]) ,
 $\|V^{\lozenge}-V\|_{1+\sigma}\leq C(\|\zeta\|_{\sigma}+\|f\|_{0})\leq$ $C\|v-v^{\lozenge}\|_{\sigma}\leq C\|v-v^{\lozenge}\|_{0}^{1-\sigma}\|v-v^{\lozenge}\|_{1}^{\sigma}$
@@
 $\|\mathcal{E}_{14}\|_{1}$ $=\|V-V^{\lozenge}\|_{1}\leq\|V-V^{\lozenge}\|_{1+\sigma}\leq C\delta^{\frac{1}{2}}\lambda^{-\tau(1-\sigma)}\leq C\delta\lambda.$
-It is remained to estimates $\mathcal{E}_{15}$. From (3.26), for $j=0,1$, we can derive
+It remains to estimates $\mathcal{E}_{15}$. From (3.26), for $j=0,1$, we can derive
 $\|\mathcal{E}_{15}\|_{j}\leq C\|v-v^{\lozenge}\|_{j}\leq C\delta^{\frac{1}{2}}\lambda^{\tau j-\tau}\leq\delta\lambda^{1-\tau(1-j)},$
@@
 (3.35) $\|\mathcal{E}_{1}\|_{0}\leq C\delta\lambda^{1-\tau},\quad\|\mathcal{E}_{1}\|_{1}\leq C\delta\lambda.$
-Step 3.Adding the second rank-one matrix.
-We also mollify $(v^{\lozenge},w^{\lozenge})$ to get $(\tilde{v}^{\lozenge},\tilde{w}^{\lozenge})$ at length-scale $\lambda^{1-2\tau}$.
-Then by (3.28) and Lemma A.1, one also has
+[…]
 (3.36) $\begin{split}&\|\tilde{v}^{\lozenge}-v^{\lozenge}\|_{1}+\|\tilde{w}^{\lozenge}-w^{\lozenge}\|_{1}\leq C\delta^{1/2}\lambda^{1-\tau},\\ &\|\tilde{v}^{\lozenge}\|_{2+j}+\|\tilde{w}^{\lozenge}\|_{2+j}\leq C\delta^{1/2}\lambda^{\tau+(2\tau-1)j},\end{split}$
@@
 Then (3.10) follows from (3.21) and
-dist(supp $\textit{dist}(\textrm{supp }a,\textrm{supp }\rho)=\lambda^{-\tau}.$
-Furthermore, same as Step 2, using (3.16),(3.15) and (3.28), one is able to derive
+$\textit{dist}(\textrm{supp }a,\textrm{supp }\rho)=\lambda^{-\tau}.$
+Furthermore, similar to Step 2, using (3.16),(3.15) and (3.28), one is able to derive
 (3.37) $\begin{split}\|v^{\star}-v^{\lozenge}\|_{0}\leq&\lambda^{1-2\tau}\|\Gamma_{1}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{0}\leq C\delta\lambda^{1-2\tau},\\ \|v^{\star}-v^{\lozenge}\|_{1}\leq&\lambda^{1-2\tau}\|\Gamma_{1}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{1}\leq C\delta,\\ \|v^{\star}-v^{\lozenge}\|_{2}\leq&\lambda^{1-2\tau}\|\Gamma_{1}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{2}\leq C\delta\lambda^{2\tau-1}.\end{split}$
@@
 (3.38) $\begin{split}\|w^{\star}-w^{\lozenge}\|_{j}\leq&C(j)\lambda^{1-2\tau}(\|\Gamma_{1}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{j}\|\tilde{v}^{\lozenge}\|_{1}\\ &\quad+\|\Gamma_{1}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{0}\|\tilde{v}^{\lozenge}\|_{j+1}\\ &\quad+\|\Gamma_{2}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{j}\|\nabla\Phi_{2}\|_{0}\\ &\quad+\|\Gamma_{2}(a(\cdot),\lambda^{2\tau-1}\Phi_{2}(\cdot))\|_{0}\|\nabla\Phi_{2}\|_{j})\\ \leq&C(j)\delta^{1/2}\lambda^{(2\tau-1)(j-1)},\,\,j=0,1,2.\end{split}$
-So combining with (3.26), (3.27), (3.37) and (3.38), we have
+By (3.26), (3.27), (3.37) and (3.38), we have
 $\|v^{\star}-v\|_{0}+\|w^{\star}-w\|_{0}\leq$ $\|v^{\star}-v^{\lozenge}\|_{0}+\|v^{\lozenge}-v\|_{0}+\|w^{\star}-w^{\lozenge}\|_{0}+\|w^{\lozenge}-w\|_{0}$
@@
 $\leq$ $C(\delta^{1/2}\lambda^{\tau}+\lambda^{2\tau-1})\leq C\delta^{1/2}\lambda^{2\tau-1},$
-and (3.6) we also gets (3.13) since $\tau>1$.
+and (3.6) we also get (3.13) since $\tau>1$.
 Finally, we shall control the second deficit. A direct calculation implies
-$\mathcal{E}_{2}:$ $=D(v^{\star},w^{\star})-(D(v^{\lozenge},w^{\lozenge})+a^{2}\nabla\Phi_{2}\otimes\nabla\Phi_{2})=\sum_{k=1}^{5}\mathcal{E}_{2k}$
+$\mathcal{E}_{2}:$ $=D(v^{\star},w^{\star})-\left[D(v^{\lozenge},w^{\lozenge})+a^{2}\nabla\Phi_{2}\otimes\nabla\Phi_{2}\right]=\sum_{k=1}^{5}\mathcal{E}_{2k}$
 with
-$\mathcal{E}_{21}:=$ $\lambda^{1-2\tau}(\partial_{t}\Gamma_{1}\partial_{s}\Gamma_{1}\textrm{sym}(\nabla\Phi_{2}\otimes\nabla a)-\Gamma_{1}\nabla^{2}\tilde{v}^{\lozenge}+\partial_{s}\Gamma_{2}\textrm{sym}(\nabla\Phi_{1}\otimes\nabla a)),$
-$\mathcal{E}_{22}:=$ ${\lambda}^{-\tau}\partial_{s}\Gamma_{1}\textrm{sym}((\nabla v^{\lozenge}-\nabla\tilde{v}^{\lozenge})\otimes\nabla a)+\partial_{t}\Gamma_{1}\textrm{sym}((\nabla v^{\lozenge}-\nabla\tilde{v}^{\lozenge})\otimes\nabla\Phi_{1}),$
+$\mathcal{E}_{21}:=$ $\lambda^{1-2\tau}\left[\partial_{t}\Gamma_{1}\partial_{s}\Gamma_{1}\textrm{sym}(\nabla\Phi_{2}\otimes\nabla a)-\Gamma_{1}\nabla^{2}\tilde{v}^{\lozenge}+\partial_{s}\Gamma_{2}\textrm{sym}(\nabla\Phi_{1}\otimes\nabla a)\right],$
+$\mathcal{E}_{22}:=$ ${\lambda}^{-\tau}\partial_{s}\Gamma_{1}\textrm{sym}\left[(\nabla v^{\lozenge}-\nabla\tilde{v}^{\lozenge})\otimes\nabla a\right]+\partial_{t}\Gamma_{1}\textrm{sym}\left[(\nabla v^{\lozenge}-\nabla\tilde{v}^{\lozenge})\otimes\nabla\Phi_{1}\right],$
 $\mathcal{E}_{23}:=$ $\tfrac{1}{2}\lambda^{-2\tau}(\partial_{s}\Gamma_{1})^{2}\nabla a\otimes\nabla a,$
 $\mathcal{E}_{24}:=$ $(V^{\star}-V^{\lozenge})\cdot\mathrm{Id},\,\,(\text{here }\Delta V^{\star}=F(v^{\star}))$
-$\mathcal{E}_{25}:=$ $((v^{\star}-v^{\lozenge})\cot\Theta)\cdot\mathrm{Id}.$
+$\mathcal{E}_{25}:=$ $\left[(v^{\star}-v^{\lozenge})\cot\Theta\right]\cdot\mathrm{Id}.$
 Repeating the estimating procedure in Step 2, we are able to get
@@
 In fact, when bounding the $C^{1}$ norm of the second matrix error, we only need to keep in mind that $\lambda^{2\tau-1}$(and $\lambda^{\tau}$) plays the same role as $\lambda^{\tau}$(and $\lambda$ repsectively) in Step 2.
-Therefore, the final matrix error can be calculated as follows
+Therefore, the final matrix error can be calculated as follows:
 $\mathcal{E}$ $=D(v^{\star},w^{\star})-(D(v,w)+\rho^{2}(\textrm{Id}+H))$
@@
 4. Proof of Theorem 1.1 (2)
-As discussed in the beginning of Section 3, we show Theorem 1.1 (2) by seeking infinitely many symmetric $2\times 2$ matrixes $A$’s and $C^{1,\alpha}(\alpha<\frac{1}{5})$ functions $v$’s such that
+As discussed at the beginning of Section 3, we show Theorem 1.1 (2) by seeking infinitely many symmetric $2\times 2$ matrixes $A$’s and $C^{1,\alpha}(\alpha<\frac{1}{5})$ functions $v$’s such that
 $\begin{cases}-\mbox{curl curl }A=1,\\ A=D(v,w)=\mathrm{sym}\nabla w+\frac{1}{2}\nabla v\otimes\nabla v-(v\cot\Theta)\cdot\mathrm{Id}+V\cdot\mathrm{Id},\end{cases}$
-where $V$ solves the following Dirichlet problem:
-(4.1) $\Delta V&=F(v)=2\nabla v\cdot\nabla\cot\Theta+v\Delta\cot\Theta,\mbox{ in }\Omega,\\ V&=0,\mbox{ on }\partial\Omega,\end{aligned}\right.$
-We will solve the first system about $A$ together with a subsolution by seeking solutions to some Possion equations and the second system by constructing a sequence of adapted subsolutions via convex integration.
+where $V$ satisfies the following Dirichlet problem:
+(4.1) $\left\{\begin{aligned} \Delta V&=F(v)=2\nabla v\cdot\nabla\cot\Theta+v\Delta\cot\Theta,\mbox{ in }\Omega,\\ V&=0,\mbox{ on }\partial\Omega,\end{aligned}\right$.
+We will solve the first system about $A$ together with a subsolution by seeking solutions to some Poisson equations and the second system by constructing a sequence of adapted subsolutions via convex integration.
 Proof of Theorem 1.1 (2).
 We separate the proof into four steps.
-Step 1. Matrix and initial subsolutions. We construct matrix $A$ and initial subsolutions at the same time. Let $u$ be the solution to the following Dirichlet problem,
-$\Delta u&=0,\mbox{ in }\Omega,\\ u&=g,\mbox{ on }\partial\Omega,\end{aligned}\right.$
+Step 1. Matrix and initial subsolutions. We construct the matrix $A$ and initial subsolutions at the same time. Let $u$ be the solution to the following Dirichlet problem,
+$\left\{\begin{aligned} \Delta u&=0,\mbox{ in }\Omega,\\ u&=g,\mbox{ on }\partial\Omega,\end{aligned}\right$.
 and $\psi$ be the solution of the following Dirichlet problem,
-$-\Delta\psi&=1-\det\nabla^{2}u,\mbox{ in }\Omega,\\ \psi&=0,\mbox{ on }\partial\Omega.\end{aligned}\right.$
+$\left\{\begin{aligned} -\Delta\psi&=1-\det\nabla^{2}u,\mbox{ in }\Omega,\\ \psi&=0,\mbox{ on }\partial\Omega.\end{aligned}\right$.
 Define
@@
 where $U$ is the solution to
-$\Delta U&=F(u):=2\nabla u\cdot\nabla\cot\Theta+u\Delta\cot\Theta,\mbox{ in }\Omega,\\ U&=0,\mbox{ on }\partial\Omega.\end{aligned}\right.$
+$\left\{\begin{aligned} \Delta U&=F(u):=2\nabla u\cdot\nabla\cot\Theta+u\Delta\cot\Theta,\mbox{ in }\Omega,\\ U&=0,\mbox{ on }\partial\Omega.\end{aligned}\right$.
 Note that $|\sin\Theta|\geq c_{2}>0$ as in the assumptions of Theorem 1.1.
@@
 $=1+\cot\Theta\Delta u=1.$
-With $\Delta u=0$, we know $01-\det\nabla^{2}u=1+(\partial^{2}_{11}u)^{2}+(\partial^{2}_{12}u)^{2}>$ 0, and hence by the strong maximum principle, we have
+With $\Delta u=0$, we know $1-\det\nabla^{2}u=1+(\partial^{2}_{11}u)^{2}+(\partial^{2}_{12}u)^{2}>0$, and hence by the strong maximum principle, we have
 $\psi(x)>0,\mbox{ for any }x\in\Omega.$
@@
 where $\rho_{0}=\sqrt{\psi}$ and $H_{0}$ is the zero matrix. Moreover, $v_{0}|_{\partial\Omega}=u|_{\partial\Omega}=g$.
-Note that $(v_{0},w_{0})$ and $A$ are not unique. For example, taking $v_{0}$ as before and $w_{0}\in\operatorname{\mathbb{R}}^{2}$ as arbitrary constant vector will also work.
-Step 2. Iteration. With Proposition 3.1, for any $q\in\operatorname{\mathbb{N}}$, we construct inductively $(v_{q},w_{q},\rho_{q},H_{q})$ such that
+Note that $(v_{0},w_{0})$ and $A$ are not unique. For example, taking $v_{0}$ as before and $w_{0}\in\operatorname{\mathbb{R}}^{2}$ as an arbitrary constant vector will also work.
+Step 2. Iteration. We define the amplitude sequence $\delta_{q}$ and the frequency sequence $\lambda_{q}$ inductively by
+$\delta_{1}=\max_{x\in\overline{\Omega}}\rho_{0}^{2},\quad\lambda_{q+1}=\lambda_{q}^{b},\quad\lambda_{q}=M\delta_{q}^{-\frac{1}{2\beta}},$
+where $\beta$ is a arbitrary fixed constant in $(0,\frac{1}{5})$, and
+(4.2) $b=1+\frac{9\sigma}{1-5\beta}\quad(\text{then }5\beta(b-1)+9\sigma=b-1),$
+and $M$ is a large constant to be fixed. We then decomposes $\overline{\Omega}$ according to the level sets of $\rho_{0}$,
+$\Omega_{q}=\left\{x\in\overline{\Omega}:\,\rho_{0}(x)\geq 2\delta_{q+2}^{1/2}\right\},\,\,\tilde{\Omega}_{q}=\left\{x\in\overline{\Omega}:\,\rho_{0}(x)\geq\frac{3}{2}\delta_{q+2}^{1/2}\right\}$
+for any $q\in\mathbb{N}$. With Proposition 3.1, for any $q\in\operatorname{\mathbb{N}}$, we inductively construct $(v_{q},w_{q},\rho_{q},H_{q})$ such that the following hold:
   • $(1)_{q}$
@@
   • $(2)_{q}$
-    If $x\in\overline{\Omega}\setminus\tilde{\Omega}_{q}$, then $(v_{q},w_{q},\rho_{q},H_{q})(x)=(v,w,\rho,H)(x).$
+    If $x\in\overline{\Omega}\setminus\tilde{\Omega}_{q}$, then $(v_{q},w_{q},\rho_{q},H_{q})(x)=(v_{0},w_{0},\rho_{0},H_{0})(x).$
   • $(3)_{q}$
@@
   • $(4)_{q}$
-    If $x\in\tilde{\Omega}_{j}$ for some $j\geq q,$ we get the following estimates
+    If $x\in\tilde{\Omega}_{j}$ for some $j\geq q,$ we get the following estimates:
     $\rho_{q}\geq\frac{3}{2}\delta_{j+2}^{1/2},\quad|\nabla\rho_{q}|\leq\delta_{j+1}^{1/2}\lambda_{j+2},\quad\frac{|\nabla\rho_{q}|}{\rho_{q}}\leq\lambda_{j+2},$
@@
     where $\overline{C}$ is the constant in the conclusion of Proposition 3.1.
-Here the amplitude sequence $\delta_{q}$ and the frequency sequence $\lambda_{q}$ are inductively defined by
-$\delta_{1}=\max_{x\in\overline{\Omega}}\rho^{2},\quad\lambda_{q+1}=\lambda_{q}^{b},\quad\lambda_{q}=M\delta_{q}^{-\frac{1}{2\beta}},$
-where arbitrary constant $\beta\in(0,\frac{1}{5}),$
-(4.3) $b=1+\frac{9\sigma}{1-5\beta}\quad(\text{then }5\beta(b-1)+9\sigma=b-1).$
-and $M$ is a large constant to be fixed.
 Indeed, $(v_{0},w_{0},\rho_{0},H_{0})$ has been constructed in Step 1 and it is not hard to verify $(1)_{0}-(4)_{0}$.
-For $q\geq 1,$ using Proposition 3.1 and following Step 2 in the proof of Proposition 2.1 of [4], we are able to inductively get $(v_{q},w_{q},\rho_{q},H_{q}).$ Since the proof is exactly same, we omit the details.
+Similar to Step 2 of the proof of Proposition 2.1 in [Cao], by using Proposition 3.1 we are able to inductively get $(v_{q},w_{q},\rho_{q},H_{q})$ for $q\geq 1$.
+Since the proof is exactly the same as in [Cao], we omit the details.
 Step 3.
-Taking limit.
+Taking limits.
 By $(3)_{q}$ and $(5)_{q}$, we know $(\rho_{q},H_{q})$ is a Cauchy sequence in $C^{0}(\overline{\Omega})$, and $(v_{q},w_{q})$ is a Cauchy sequence in $C^{1}(\overline{\Omega})$, respectively.
@@
 for some functions $\bar{v}$ and $\bar{w}$.
-Moreover, for any $\alpha\in(0,\frac{1}{5})$ and $\varepsilon>0$, we can choose $\beta\in(\alpha,\frac{1}{5})$ in Step 2, by $(4)_{q}-(5)_{q}$, $(4)_{q+1}$ and interpolation, we have for any $\alpha\in(0,\frac{\beta}{b^{2}})$,
+Moreover, for any $\alpha\in(0,\frac{1}{5})$ and $\varepsilon>0$, we can choose $\beta\in(\alpha,\frac{1}{5})$ in Step 2.
+By $(4)_{q}-(5)_{q}$, $(4)_{q+1}$ and interpolation, we have for any $\alpha\in(0,\frac{\beta}{b^{2}})$,
 $\|v_{q}-v_{q+1}\|_{1+\alpha}+\|w_{q}-w_{q+1}\|_{1+\alpha}\leq 4\delta_{q+1}^{(1-\alpha)/2}\delta_{q+2}^{\alpha/2}\lambda_{q+3}^{\alpha}\leq M^{\beta}\lambda_{q+1}^{\alpha b^{2}-\beta},$
 which implies $(v_{q},w_{q})$ uniformly converges to $(\bar{v},\bar{w})$ in $C^{1,\alpha}(\overline{\Omega})$.
-Since $\alpha$ and $b$ can be taken arbitrarily close to 00 and $1$ by (4.3), respectively, $(\bar{v},\bar{w})\in C^{1,\alpha}({\Omega}),$ for any $\alpha<\beta$.
+Since $\alpha$ and $b$ can be taken arbitrarily close to 0 and $1$ by (4.2), respectively, $(\bar{v},\bar{w})\in C^{1,\alpha}({\Omega}),$ for any $\alpha<\beta$.
 From $(5)_{q}$, for any $\epsilon>0,$ we also have
@@
 Step 4.
-Very weak solution.
+Very weak solutions.
 We will show $\bar{v}$ is a very weak solution to (1.1).
 On the one hand, by construction, it is easy to see $(\bar{v},\bar{w})|_{\partial\Omega}=(v_{0},w_{0})|_{\partial\Omega}$.
-On the other hand, $\rho_{q}\to 0$ implies $A-D(\bar{v},\bar{w})=0.$ Thus for any $\phi\in C^{\infty}_{0}(\Omega)$, by integration by parts and curl crul $\mbox{curl crul }A=-1$, we have
-$-\int_{\Omega}\phi\sin\Theta\mathrm{d}x=$ $\int_{\Omega}\mathrm{Trace}(H_{\phi\sin\Theta}\cdot A)\mathrm{d}x$
-$=$ $\int_{\Omega}\mathrm{Trace}(H_{\phi\sin\Theta}\cdot\frac{1}{2}\nabla\bar{v}\otimes\nabla\bar{v})\mathrm{d}x-\int_{\Omega}\bar{v}\Delta(\phi\cos\Theta)\mathrm{d}x,$
-where we have used curl $\mbox{curl curl}(\mathrm{sym}\nabla\bar{w})=0$ and notation
+On the other hand, $\rho_{q}\to 0$ implies $A-D(\bar{v},\bar{w})=0.$ Thus for any $\phi\in C^{\infty}_{0}(\Omega)$, by integration by parts and $\mbox{curl curl }A=-1$, we have
+$-\int_{\Omega}\phi\sin\Theta dx=$ $\int_{\Omega}\mathrm{Trace}(H_{\phi\sin\Theta}\cdot A)dx$
+$=$ $\int_{\Omega}\mathrm{Trace}(H_{\phi\sin\Theta}\cdot\frac{1}{2}\nabla\bar{v}\otimes\nabla\bar{v})dx-\int_{\Omega}\bar{v}\Delta(\phi\cos\Theta)dx,$
+where we have used $\mbox{curl curl}(\mathrm{sym}\nabla\bar{w})=0$ and the notation
 $H_{f}=\left(\begin{array}[]{rr}\partial_{22}^{2}f&-\partial_{12}^{2}f\\ -\partial_{12}^{2}f&\partial_{11}^{2}f\\ \end{array}\right)\text{ for any $C^{2}$ function }f.$
 Then it implies
-$0=\int_{\Omega}\partial_{1}\bar{v}\partial_{2}\bar{v}\partial_{12}^{2}(\phi\sin\Theta)$ $-\frac{1}{2}(\partial_{1}\bar{v})^{2}\partial_{22}^{2}(\phi\sin\Theta)-\frac{1}{2}(\partial_{2}\bar{v})^{2}\partial_{11}^{2}(\phi\sin\Theta)\mathrm{d}x$
-$+\int_{\Omega}\bar{v}\Delta(\phi\cos\Theta)\mathrm{d}x-\int_{\Omega}\phi\sin\Theta\mathrm{d}x.$
+$0=\int_{\Omega}\partial_{1}\bar{v}\partial_{2}\bar{v}\partial_{12}^{2}(\phi\sin\Theta)$ $-\frac{1}{2}(\partial_{1}\bar{v})^{2}\partial_{22}^{2}(\phi\sin\Theta)-\frac{1}{2}(\partial_{2}\bar{v})^{2}\partial_{11}^{2}(\phi\sin\Theta)dx$
+$+\int_{\Omega}\bar{v}\Delta(\phi\cos\Theta)dx-\int_{\Omega}\phi\sin\Theta dx.$
 Finally, since $\epsilon$ can be arbitrary and the initial subsolution is also not unique (note each $(v_{q},w_{q})$ can be an initial subsolution), (1.1) admits infinitely many very weak solutions $\bar{v}$’s.
@@
 Appendix A Useful lemmas and proofs of Lemmas 3.2-3.3
-In this appendix, we provide the sketch proofs of Lemmas 3.2-3.3 for completeness, which are as same as those in the first author [4]. We begin with the following two useful lemmas.
-Lemma A.1 ([9]).
+In this appendix, we provide the sketch proofs of Lemmas 3.2-3.3 for completeness, which are as same as those in [Cao]. We begin with the following two useful lemmas.
+Lemma A.1 ([CDS12]).
 For any $r,s\geq 0$ and $0<\alpha\leq 1$, we have
@@
 $\|(h_{1}h_{2})*\phi_{l}-(h_{1}*\phi_{l})(h_{2}*\phi_{l})\|_{r}\leq Cl^{2\alpha-r}\|h_{1}\|_{\alpha}\|h_{2}\|_{\alpha},$
-with constant $C$ depending only on $s,r,\alpha,\phi$.
-Lemma A.2 ([4]).
-Let $\Sigma\subset\operatorname{\mathbb{R}}^{2}$ be a simply connected open bounded set with smooth boundary and $H:\Sigma\to\operatorname{\mathbb{R}}^{2\times 2}$ a smooth $2\times 2$ symmetric positive definite matrix valued function, such that for some $10<\alpha<$ 1, and $c,M\geq 1$,
+with the constant $C$ depending only on $s,r,\alpha,\phi$.
+Lemma A.2 ([Cao]).
+Let $\Sigma\subset\operatorname{\mathbb{R}}^{2}$ be a simply connected open bounded set with a smooth boundary and $H:\Sigma\to\operatorname{\mathbb{R}}^{2\times 2}$ a smooth $2\times 2$ symmetric positive definite matrix-valued function, such that for some $0<\alpha<1$, and $c,M\geq 1$,
 $c^{-1}\mathrm{Id}\leq H\leq c\mathrm{Id},\,\,\|H\|_{C^{\alpha}(\Sigma)}\leq M.$
-Then there exists a smooth diffeomorphism $\Phi:\Sigma\to\operatorname{\mathbb{R}}^{2}$ and a smooth positive function $\vartheta:\Sigma\to\operatorname{\mathbb{R}}$ satisfying
+Then there exist a smooth diffeomorphism $\Phi:\Sigma\to\operatorname{\mathbb{R}}^{2}$ and a smooth positive function $\vartheta:\Sigma\to\operatorname{\mathbb{R}}$ satisfying
 $H=\vartheta^{2}(\nabla\Phi_{1}\otimes\nabla\Phi_{1}+\nabla\Phi_{2}\otimes\nabla\Phi_{2}).$
@@
 where the constants $c_{0}>0$, $C(j)\geq 1$ depend only on $\alpha,c,M$ and on $\Sigma$.
-Sketch proof of Lemma 3.2.
-Recall the definition of $\Gamma_{i}:\operatorname{\mathbb{R}}\times\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}}$, $i=1,2$,
+A sketch proof of Lemma 3.2.
+Recall the definition of $\Gamma_{i}:\operatorname{\mathbb{R}}\times\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}}$, $i=1,2$:
 $\Gamma_{1}(s,t)=\frac{s}{\pi}\sin(2\pi t),\quad\Gamma_{2}(s,t)=-\frac{s^{2}}{4\pi}\sin(4\pi t).$
-For any nonnegative integer $k$, it is not hard to show the following estimates.
+For any nonnegative integer $k$, it is not hard to show the following estimates:
 (A.1) $\begin{split}|\partial_{t}^{k}\Gamma_{1}(s,t)|+|\partial_{s}\partial^{k}_{t}\Gamma_{2}(s,t)|&\leq C|s|,\\ |\partial_{s}\partial_{t}^{k}\Gamma_{1}(s,t)|&\leq C,\\ |\partial_{t}^{k}\Gamma_{2}(s,t)|&\leq Cs^{2}.\end{split}$
-The derivatives of $\Gamma_{i}$, $i=1,2$ could be calculated directly by the chain rule, then the desired estimates are directly following by the assumptions on $\rho,H$, Lemma A.2 and (A.1). ∎
-Sketch proof of Lemma 3.3.
+The derivatives of $\Gamma_{i}$, $i=1,2$ could be calculated directly by the chain rule, then the desired estimates directly follow from the assumptions on $\rho,H$, Lemma A.2 and (A.1). ∎
+A sketch proof of Lemma 3.3.
 By (3.15), (3.16), (3.17), (3.18), (3.24) and (3.25), we get